### v9

# Balancing LoRA Performance and Efficiency with Simple Shard Sharing

Abstract
Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA), effectively reduce the number of trainable parameters in Large Language Models (LLMs). However, as model scales continue to grow, the demand for computational resources remains a significant challenge. Existing LoRA variants often struggle to strike an optimal balance between adaptability (model performance and convergence speed) and efficiency (computational overhead, memory usage, and initialization time). This paper introduces Fossil(Framework for Optimal Shard Sharing Integration in LoRA), a novel PEFT approach that addresses this trade-off through a simple shard-sharing mechanism. Fossil leverages the insight that a low-rank adaptation can be achieved by decomposing the weight matrix into multiple fragment matrices and utilizing a shared, trainable common fragment. This method constructs the low-rank update matrix through the replication of these shared, partitioned shards. We also propose a hardware-efficient and broadly applicable implementation for Fossil. Extensive experiments conducted on a range of tasks, alongside a systematic analysis of computational performance, demonstrate Fossil’s superiority. The results show that Fossil significantly outperforms standard LoRA and its prominent variants in both model performance metrics and computational efficiency, including initialization speed and training throughput. By effectively balancing expressive power and resource utilization, Fossil offers a compelling solution for efficiently adapting large-scale models.
[Uncaptioned image]https://github.com/huggingface/peft
1 Introduction
Fine-tuning Large Language Models (LLMs) [16, 17] is a prevalent methodology for adapting these models to specific downstream tasks. However, the comprehensive fine-tuning of all parameters typically incurs substantial computational costs. Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques [23] have been developed to mitigate the training expenditure associated with these large-scale models. Among such techniques, Low-Rank Adaptation (LoRA) [8] has distinguished itself as one of the most prominent PEFT methods. LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead. LoRA constructs this low-rank adaptation matrix through an intuitive design, positing that the weight update $\Delta\bm{W}$ can be approximated by the product of two lower-rank matrices, $\bm{B}\bm{A}\approx\Delta\bm{W}$. Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
The improvements of LoRA are mimicked in two streams: (1) Adaptability [6, 12, 2]: This refers to the celerity with which the method converges to an optimal or near-optimal state. The approximation must exhibit a representational capacity comparable to that of the original, full $\Delta\bm{W}$. Extensive experiments have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants [7, 14, 20]. By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed. Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients. Certain solutions may improve the performance of finetuning, but makes the initialization process significantly more complex or increase the training cost, leading to a worse complexity. (2) Efficiency [11, 19, 22]: This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead. Optimizing LoRA from an efficiency perspective can lead to reduced VRAM consumption and an accelerated training process. However, further simplification of the LoRA structure may result in an additional loss in performance. However, the critical balance between achieving highly expressive architectures and maintaining algorithmic efficiency has often received insufficient attention.
In this work, we systematically investigate the Pareto frontier of the trade-off between Adaptability and Efficiency in PEFT. We contend that an ideal PEFT method must adeptly reconcile these two crucial aspects. To this end, we first conduct a series of foundational experiments focusing on these dimensions, encompassing a simulated pre-training and fine-tuning pipeline, computational complexity analysis, and initialization time tests. Our analysis reveals that while LoRA-like methods leveraging shard sharing can achieve excellent computational efficiency, their expressive power may be constrained by such sharing, sometimes leading to performance below the standard LoRA baseline. Conversely, SVD-based modifications to the original LoRA often exhibit strong adaptability, yet their practical efficiency can be undermined by complex initialization procedures and more intricate forward and update mechanisms.
Considering the aforementioned trade-off, our investigation reveals that a surprisingly simple strategy—constructing a low-rank matrix through the replication of a shared, partitioned shard—can achieve a compelling balance across these criteria. This approach, which we introduce as Fossil,  exhibits notable strengths. In terms of adaptability, Fossil demonstrates superior approximation fidelity compared to LoRA and several of its variants. Its relatively substantial gradient norm (as illustrated in Figure 4) promotes rapid convergence, a key characteristic of its high adaptability. Furthermore, concerning efficiency, Fossil benefits from rapid initialization; its hardware-friendly, repetition-based initialization facilitates swift processing, particularly when scaling to large models with substantial parameter counts. The formulation of Fossil can also be expressed in an efficient and mathematically equivalent form that exhibits favorable time and space complexity.
Table 1: A variety of LoRA variants are listed, each with its specific update formulation and initialization strategy for the low-rank matrices. The differences between these methods are compared in a clear and intuitive manner. e denotes efficient form.
Method Forward Initialization
LoRA [8] $y=\bm{W}_{0}x+\bm{BA}x$ $\bm{A}\sim N(0,\sigma^{2})$ $\bm{B}\sim 0$
PiSSA $y=\bm{W}_{0}x+\bm{BA}x$ $\bm{A}=U_{[:,:r]}S_{[:r,:r]}^{1/2},\ \bm{B}=S_{[:r,:r]}^{1/2}V_{[:,:r]}^{\top}$
AdaLoRA $y=\bm{W}^{(0)}x+\bm{P\Lambda Q}x$ $\bm{\Lambda}\sim 0,\ \bm{P},\bm{Q}\sim N(0,\sigma^{2})$
DoRA $y=\bm{m}(\ \bm{W}_{0}x+\bm{BA}x~{}/~{}{\|\bm{W}_{0}+\bm{BA}\|_{c}})$ $\bm{A}\sim\mathrm{Rect.KaimingUnif},\ \bm{B}\sim 0$
ProLoRA $y=\bm{W}_{0}x+(\bm{B_{u}}\oplus_{h}\dots)$ $(\bm{A_{u}}\oplus_{v}\dots)x$ $\bm{A_{u}}\sim\mathrm{KaimingUnif},\ \bm{B_{u}}\sim 0$
MoS $y=\bm{W}_{0}x+\bm{B}^{s}\bm{A}^{s}x$ $\bm{A}^{\text{pub/pri}},\bm{B}^{\text{pub/pri}}\sim 0$
Fossil(Ours) $y=\bm{W_{0}}x+\operatorname{expand}(\bm{D})x$ $\bm{D}\sim 0$
Fossile(Ours) $y=\bm{W_{0}}x+\bm{D^{\top}S}$ $\bm{D}\sim 0$
We perform scaling experiments to verify our preliminary conclusions. Our extensive evaluation shows that Fossil demonstrates excellent performance on understanding and generation tasks, significantly surpassing other baseline method. Moreover, we present the training dynamics i.e., loss curves, which Fossil also showcase a significantly faster convergence rate compared to other baselines. In generation tasks, Fossil demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variants across the board. This demonstrates the feasibility of the framework. Furthermore, we validate the efficiency of Fossil  which enables it to significantly outperform LoRA and its variants in both computational efficiency and memory usage.
Our contributions can be summarized as threefolds:
  1. 1.
    In Section 3, we discuss the balancing Between Adaptability and Efficiency in LoRA-like methods with highly artificial controlled dataset to make the preliminary verification transparently and easy for replication.
  2. 2.
    In Section 4, we proposed Fossil, an efficient and adaptable structure with shard sharing mechanism.
  3. 3.
    In Section 5, we scale the experiments across a wide range of datasets and models, validated its effectiveness through extensive experiments.
def init(in_feats: int, out_feats: int):
weight = nn.Parameter(
torch.empty((out_features, in_features)) )
D = nn.Parameter(torch.zeros(self.r, out_features))
def forward(self, x):
result = F.linear(x, self.weight) # x: [B, T, C]
y = result + torch.sum(x.reshape
(*x.shape[:-1], x.size(-1)//self.r, self.r)\\
, dim=-2) @ self.fossil
return y
Figure 1: Left. Schematic diagram of the forward computation and initialization methods in LoRA and Fossil. Right. PyTorch-style pseudocode illustrating the implementation of Fossil.
2 Preliminaries and Related Works
Low-Rank Adaptation (LoRA).
Parameter-Efficient Fine-Tuning (PEFT) refers to a family of techniques designed to adapt large pre-trained models to downstream tasks while minimizing the number of trainable parameters, thereby reducing computational and memory overhead. Among diverse methods, Low-Rank Adaptation (LoRA) has gained significant prominence. It operates on the principle that the change in weights during model adaptation often possesses a low intrinsic rank. Instead of fine-tuning the entire pre-trained weight matrix $\bm{W}_{0}\in\mathbb{R}^{d\times k}$, LoRA introduces a low-rank decomposition to represent the update. Consider a simple linear projection with input $x\in\mathbb{R}^{d}$ and output $y\in\mathbb{R}^{k}$, LoRA adapts the following forward pass:
$y=(\bm{W}_{0}+\bm{\Delta W})x\approx\bm{W}_{0}x+\bm{BA}x,\ \text{where}\ \bm{B}\in\mathbb{R}^{d\times r},\ \bm{A}\in\mathbb{R}^{r\times k}.$ (1)
Here, $\bm{A}$ and $\bm{B}$ are low-rank matrices, with the rank $r$ being significantly smaller than the original dimensions i.e., $r\ll\min(d,k)$. During the fine-tuning process, the original weights $\bm{W}_{0}$ are kept frozen, and only the parameters within matrices $\bm{A}$ and $\bm{B}$ are trained. Specifically, LoRA initializes $\bm{A}$ with Gaussian noise $A\sim N(0,\sigma^{2})$ with small $\sigma$ and $\bm{B}$ with zeros, ensuring that $\bm{BA}=0$ at the start, preserving the pre-trained model’s output.
Improvements of LoRA.
LoRA is the low rank adaptation towards full-param finetuning, and intuitively it downperforms than it. Several works propose diverse methods towards a better convergence and adaptability of LoRA. One compelling venue is to change the form of LoRA. PiSSA [14] optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), OLoRA [3] leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $\bm{A}$ and $\bm{B}$ by computing the initial gradient, thereby closely approximating full fine-tuning. LoRA+ extended this method by introducing independent learning rates for matrices $\bm{A}$ and $\bm{B}$ with a fixed ratio, improving the method’s efficiency. DoRA [12] decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately. This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning. The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential. However, they also introduce more complex initialization steps and increase preprocessing time.
3 No Free Launch: Balancing Between Adaptability and Efficiency
This section elucidates the fundamental trade-off inherent in LoRA-style PEFT techniques: the delicate balance between their adaptability and efficiency. Adaptability, in this context, refers to the capacity of a given method to emulate the performance benchmarks set by full-parameter fine-tuning. Conversely, efficiency encompasses the method’s judicious use of computational resources, specifically time and memory. We utilize highly artificial controlled dataset and model with relatively small parameter count to make the verification transparently and easy for replication.
We considered diverse methods  [^1] We have not included methods such as LoRA-GA [21] or LoRA+ [7] in our current analysis. While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider Fossil to be largely orthogonal to them. Consequently, the analytical techniques employed in their study may still offer valuable insights for Fossil.: (1) Full-parameter finetuning [13]. (2) LoRA [8]. (3) Alternatives to LoRA w/ different architectures, including: PiSSA [14], VeRA [11], DoRA [12] and MoRA [10]. (4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA [19], MoS [22]. (1) An overview of their forward form, initialization method can be found at Table 1.
3.1 Empirically Benchmarking the Adaptability of LoRA Variants
Experimental Setup.
Parameter-efficient adaptation methods, particularly those leveraging low-rank principles, typically constrain trainable parameters by applying low-rank decompositions either to newly introduced adapter matrices or to the updates of pre-existing model weights. To rigorously evaluate such strategies, we selected a deliberately minimalistic base model: a single-layer MLP designed to process a series of features and yield outputs. This model is initially pre-trained to fit some sinusoidal functions using a constrained set of data points. Following this pre-training, the target function is subtly altered, and an additional dataset sampled from this modified function is employed for training to assess the adaptation performance of various fine-tuning techniques. Comprehensive details regarding the experimental settings are elaborated in Appendix B.
Results.
Figure 2 illustrates the comparative adaptability of different methods. We utilize the minimum validation loss achieved by each approach as an indicator of its expressive capacity when approximating the performance of full-parameter fine-tuning. The results clearly demonstrate that methods leveraging singular value decomposition (SVD), such as PiSSA, attain a relatively low loss. Conversely, efficiency-focused techniques like MoS exhibit higher losses. A plausible explanation for this discrepancy is that such methods further decompose LoRA matrices into shared components which may inherently constrain their expressive power. Our method Fossil reaches a relatively advanced performance comparing to other variants.
Figure 2: No Free Launch Experiment. Left. The training loss curves of all methods. Middle. Initialization time w/ parameters. Right. Training time w/ parameters.
3.2 Efficiency Analysis of LoRA Variants
Metrics.
We evaluate the efficiency of LoRA-like variants from two primary perspectives: (1) Space and Time Complexity in Training. Space and time complexity during training are generally considered crucial criteria for evaluating PEFT methods. To benchmark these aspects, we employ the model architecture detailed in Section 3.1. We also test the real cost in our experiment section i.e., Section 5.5. (2) Initialization. Initialization time is often overlooked in theoretical complexity analyses. This oversight typically stems from the assumption that common initialization techniques (e.g., Kaiming Initialization) are computationally inexpensive and represent a one-time cost within the entire training pipeline. However, several recent advancements in LoRA and its variants incorporate matrix operations (e.g., Singular Value Decomposition - SVD) that are not inherently hardware-friendly and can pose challenges for efficient optimization and computation. Consequently, we explicitly include initialization time as a distinct evaluation metric in our experimental framework. We then progressively scale the trainable parameter count of various approaches to meticulously measure their respective time and space costs.
Results.
The efficacy (See Figure 2) of Fossil is evident: its strategic combination of parameter sharing and an efficient computational design culminates in rapid, scalable performance across both initialization and training stages. In contrast, while techniques like PiSSA demonstrate commendable adaptability, as shown in prior experiments, their reliance on computationally intensive Singular Value Decomposition for initialization significantly hampers their overall speed. Other approaches, such as VeRA and AdaLoRA, offer efficient initialization and computation; however, as previously discussed, they often achieve this at the cost of comparatively reduced adaptability.
4 Fossil: Shard Sharing for the Performance and Efficiency Tradeoff
4.1 Method Overview
In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{AB}^{T}$, where $\bm{A}\in\mathbb{R}^{d\times r}$, $\bm{B}\in\mathbb{R}^{k\times r}$, and the rank $r\ll\min(d,k)$. This approach achieves efficiency by limiting the number of parameters. However, we observe that a repeating matrix—where a small matrix is replicated to form a larger one—can also be viewed as a low-rank structure. For instance, if a matrix’s rows or shards are constructed by repeating a limited set of independent elements, its effective rank is often much smaller than its full dimensions.
Based on this insight, we propose Fossil, which defines the weight update $\Delta\bm{W}$ as a large matrix generated from a small trainable matrix $\bm{D}$ through an expansion operation. The updating of $\bm{W}$ and the forward pass can be expressed as:
$\bm{W}=\bm{W_{0}}+\Delta\bm{W}=\bm{W_{0}}+\operatorname{expand}(\bm{D}),\ \ \ y=\bm{W_{0}}x+\operatorname{expand}(\bm{D})x.$ (2)
Here, $x\in\mathbb{R}^{b\times l\times k}$, $y\in\mathbb{R}^{b\times l\times d}$, $\bm{W_{0}}\in\mathbb{R}^{d\times k}$ is the pre-trained weight matrix, $\bm{D}\in\mathbb{R}^{r_{1}\times r_{2}}$ is a small trainable matrix with $(r_{1},r_{2})\ll\min(d,k)$, and $\operatorname{expand}(\bm{D})$ is a function that extends $\bm{D}$ to $\mathbb{R}^{d\times k}$. This structure inherently exhibits low-rank properties. Since the rows within each shard are identical, the rank of $\operatorname{expand}(\bm{D})$ is at most $N$. When $N\ll d$, $\Delta W$ is a low-rank matrix, reducing the parameter count from $d\times k$ to $N\times k$.
Regarding the expansion method, we partition the output dimension $d$ of $\bm{W_{0}}$ into $N$ shards of sizes $\{s_{1},s_{2},\dots,s_{N}\}$, where $\sum_{i=1}^{N}s_{i}=d$. Let $\bm{D}\in\mathbb{R}^{N\times k}$, where $N$ is the number of shards. For each shard $i$, its update is determined by the $i$-th row of $\bm{D}$, denoted $\bm{D_{i}}\in\mathbb{R}^{1\times k}$, repeated $s_{i}$ times to form the shard’s update matrix. Formally:
$(\operatorname{expand}(\bm{D}))^{\top}=\begin{bmatrix}(\mathbf{1}_{s_{1}}\bm{D_{1}})^{\top}\ (\mathbf{1}_{s_{2}}\bm{D_{2}})^{\top}\ \dots\ (\mathbf{1}_{s_{N}}\bm{D_{N}})^{\top}\end{bmatrix}$ (3)
Here, $\bm{1}_{s_{i}}\in\mathbb{R}^{s_{i}\times 1}$ is an all-ones vector, and $\bm{1}_{s_{i}}\bm{D_{i}}$ denotes $\bm{D_{i}}$ repeated $s_{i}$ times vertically. The shards are vertically concatenated to match the dimensions of $\bm{W_{0}}$.
4.2 Efficient Implementation of Fossil
The above formulation is effective in the initialization process, as it only needs to initialize a small $\bm{D}$. However, directly computing $\operatorname{expand}(\bm{D})x$ has a time complexity of $O(bldk)$ and memory complexity of $O(dk)$ , which can be computationally intensive. It is obvious that Fossil can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{r\times d}$, where $r$ is a tunable rank parameter. Instead of partitioning the output dimension $d$, we divide the input dimension $k$ into $r$ blocks, each of size $g=\lfloor k/r\rfloor$ (for simplicity, assume $k$ is divisible by $r$). For an input $\bm{x}\in\mathbb{R}^{b\times l\times k}$, partition it along the $k$-dimension, and sum each block along the $k$-dimension:
$\bm{x}$ $=[\bm{x^{(1)}},\bm{x^{(2)}},\dots,\bm{x^{(r)}}],\quad\bm{x^{(i)}}\in\mathbb{R}^{b\times l\times g},$ (4)
$\bm{S}$ $=\left[\sum_{j=1}^{g}\bm{x^{(1)}}_{[:,:,j]},\sum_{j=1}^{g}\bm{x^{(2)}}_{[:,:,j]},\dots,\sum_{j=1}^{g}\bm{x^{(r)}}_{[:,:,j]}\right]\in\mathbb{R}^{b\times l\times r}.$ (5)
This enjoys the following updating term and forward pass:
$\Delta\bm{Wx}=\bm{D^{\top}S},\ \ \bm{y}=\bm{W_{0}x}+\bm{D^{\top}S},\ \text{where}\ \bm{D}^{\top}\in\mathbb{R}^{d\times r}.$ (6)
Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{D}^{\top}\bm{S}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=D^{T}S$. Specifically, $\operatorname{expand}(\bm{D})\in\mathbb{R}^{d\times k}$ has rows corresponding to rows of $\bm{D}$, repeated across blocks in the $k$-dimension. E.g., if $k=6$, $r=3$, and $g=2$, the $i$-th row of $\operatorname{expand}(\bm{D})$ takes values $\bm{D}_{j,i}$ in block $j=\lceil j^{\prime}/g\rceil$, where $j^{\prime}$ is the column index. This structure avoids storing the $d\times k$ matrix explicitly, requiring only $\bm{D}\in\mathbb{R}^{r\times d}$, significantly reducing memory usage.
The efficient implementation of Fossil relies on an innovative input aggregation mechanism, namely blockwise input summation. We highlight its advantages through the following steps: (1) Input Partitioning and Aggregation: The aggregation exploits local redundancy in the input, preserving critical information while reducing the computational dimensionality. (2) Fast Computation: The cost of computing the efficient form is significantly lower than the original complexity. (3) Resource Savings: Memory usage drops comparing to original form. For example, with $k=1024$ and $r=16$, memory is reduced by about 64 times. An overall analysis of space and time complexity is analyzed in Table 5.
(a) Loss-Token
(b) Loss-Time
Figure 3: Left. Fossil consumes less GPU memory and achieves higher throughput(# tokens per second). GPU memory usage refers to the amount of memory occupied by an RTX 4090 card. Right. The results of fine-tuning LLaMA2-7B on MetaMathQA.
5 Experiments
In this section, we evaluate the performance of Fossil on various benchmark datasets. Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model. Subsequently, we evaluated the Natural Language Generation (NLG [^2] Evaluations were performed with the OpenCompass repository, and the Math dataset was evaluated using a 5-shot prompt configuration.) capabilities by fine-tuning the LLM. To ensure fair comparisons, we select methods officially included in the PEFT repository to avoid discrepancies caused by inconsistent tuning implementations.
Table 2: The results of fine-tuning RoBERTa-base using Fossil and various LoRA variants were compared on a subset of the GLUE benchmark.
Method Trainable MNLI SST-2 CoLA QNLI MRPC
LoRA 0.236% 85.63±0.01 94.03±0.02 62.40±0.71 91.37±0.97 87.98±0.23
PiSSA 0.236% 85.72±0.40 93.64±0.13 67.28±0.59 91.40±0.54 88.11±0.24
Fossil 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60
Table 3: We conducted fine-tuning of large language models using Fossil and multiple LoRA variants, evaluating their performance on GSM8K, Math, HumanEval, and MBPP benchmarks. All reported results are averaged over three independent runs to ensure robustness. The first-place entry should be highlighted in bold, and the second-place entry should be underlined.
Model Strategy Trainable GSM8K [5] Math [25] HumanEval [4] Mbpp [1]
Llama2-7B [18] LoRA 89.9M 40.75 5.22 17.74 35.15
DoRA 91.3M 42.93 6.51 21.95 36.53
PiSSA 89.9M 43.89 6.92 22.15 37.84
Fossil 87.0M 48.16 8.58 23.63 36.81
RWKV 6-7B [15] LoRA 88.1M 38.13 6.06 - -
PiSSA 88.1M 40.48 6.12 - -
Fossil 88.1M 41.73 6.52 - -
Mistral-7B [9] LoRA 94.4M 62.85 15.82 35.71 46.11
DoRA 95.8M 63.68 13.60 38.41 48.73
PiSSA 94.4M 67.01 18.13 41.28 51.37
Fossil 87.0M 68.92 18.85 42.07 61.33
Llama2-13B [18] LoRA 250M 56.18 12.60 31.79 37.82
DoRA 252M 61.56 13.60 33.50 39.25
PiSSA 250M 66.64 13.82 33.57 46.03
Fossil 255M 68.64 15.74 38.15 47.91
Qwen3-4B [24] LoRA 74.3M 84.38 15.20 73.27 78.32
DoRA 75.4M 85.11 21.73 74.20 78.77
PiSSA 74.3M 85.78 26.00 75.01 78.04
Fossil 70.1M 85.52 34.82 74.48 78.05
5.1 Experiments on Natural Language Understanding
We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC. Performance is evaluated on the development set using accuracy as the primary metric. The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU. Each experiment is conducted with 3 different random seeds, and the average performance is reported. As shown in Table 2, Fossil demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and PiSSA.
5.2 Experiment on Natural Language Generation
Setup.
To verify the generalizability of Fossil, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math and code. (1) Math: We trained our model on a 395k subset of MetaMathQA [25], a dataset bootstrapped from other math instruction tuning datasets like GSM8K [5] and MATH  [25], with higher complexity and diversity. (2) Code: We train our model on a 100k subset of CodeFeedback [26], a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval [4] and Mbpp [1]. The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing Fossil with LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of Fossil to 64, ensuring that Fossil uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
32641282560612182430rankgradient normftLoRAPiSSAFossil
Figure 4: Left. The loss curve for LLaMA2-7B fine-tuned on MetaMathQA shows Fossil’s superior fitting ability across architectures and parameter settings, with a notably rapid loss reduction in the first 100 steps, underscoring its effectiveness. Right. The initial gradient norm of model training.
Results.
As shown in Table 6, Fossil consistently achieves state-of-the-art performance across multiple models and evaluation metrics. Specifically, on LLaMA2-7B, Fossil surpasses PiSSA on most metrics while using fewer trainable parameters than both LoRA and DoRA, showcasing its parameter efficiency and superior performance. On RWKV6-7B, where all methods share the same number of trainable parameters, Fossil still achieves the best overall performance, indicating its strong adaptation capability to non-transformer architectures. Across all evaluated models, including Mistral-7B, LLaMA2-13B, and Qwen3-4B, Fossil maintains its leading position on key benchmarks, often achieving the highest scores with significantly fewer parameters—further highlighting its effectiveness and scalability.
5.3 Effect of Rank $r$
This subsection explores the upper limits of the Fossil structure by varying the rank $r$ in the Fossil matrix. Comparative experiments were conducted by fine-tuning LLaMA2-7B  [^3] We use LLaMA2-7B instead of LLaMA3-8B because we found that the LLaMA3 series is over-optimized on math-related tasks. on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when $r=16$, the Fossil structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is only 3.73. The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of $r$ impacts the model’s ability to understand unseen data. Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
Table 4: Comparing different values of rank $(r)$ on LLaMA2-7B with Fossil.
Model Rank Trainable GSM8K Math
Llama2-7B 16 21.7M 45.90 3.77
32 43.5M 46.18 7.43
64 87.0M 48.16 8.58
128 174.0M 53.49 10.08
Figure 5: Time and space complexity. e denotes efficient form.
Method Space Time
Full $O(dk)$ $O(bld(d+k))$
LoRA $O(dr+rk)$ $O(blr(d+k))$
Fossil $O(dr)$ $O(bldk)$
Fossile $O(dr)$ $O(blr(d+\frac{k}{r}))$
5.4 Gradient Norm Analysis
To further investigate the adaptive capabilities of Fossil, we analyzed its initial gradient norms in comparison to Full Fine-tuning (FT), LoRA, and PiSSA. The magnitude of initial gradient norms is often correlated with faster convergence and the ability of a model to effectively adapt during fine-tuning. Our comparative analysis, illustrated in Figure 4 (Right), reveals distinct behaviors among the evaluated methods. Standard LoRA consistently exhibits the lowest initial gradient norms across the tested ranks. While PiSSA’s gradient norms increase with higher ranks, they generally remain significantly lower than those achieved by full fine-tuning. Notably, Fossil demonstrates substantially larger initial gradient norms compared to both LoRA and PiSSA across various ranks. More importantly, the gradient norm profile of Fossil closely approximates that of Full Fine-tuning. This proximity to FT’s gradient characteristics suggests that Fossil is capable of inducing more significant initial updates, potentially leading to a training dynamic more akin to full fine-tuning.
5.5 Resource and efficiency
Table  5 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Fossil, and Bat on a single 4090 GPU. The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512. The results show that Fossil has the highest computational efficiency, being nearly 10% faster than LoRA while also being more memory-efficient. However, Bat incurs significantly higher memory usage due to large intermediate values and is slower in comparison. At the end of the table, we provide the actual resource costs for fine-tuning RWKV6 on the MetaMathQA dataset using 4 NVIDIA 4090 GPUs, with checkpoint techniques applied.
Table 5: Resource and efficiency
Devices Strategy Trainable Memory throughput
GPU $\times$ 1 LoRA 55.1M 12074 MB 3.62 kt/s
Fossil 55.1M 11052 MB 3.99 kt/s
GPU $\times$ 4 LoRA 55.1M 4$\times$15328 MB 15.6 kt/s
Fossil 55.1M 4$\times$15304 MB 16.0 kt/s
6 Conclusion
In this work, we proposed the Foissl framework, which divides pre-trained weights into multiple shards and updates them using a shared trainable matrix. This approach significantly reduces resource overhead and opens up new directions for PEFT techniques. Furthermore, we present a hardware-optimized version of Fossil, achieving remarkable gains in computational efficiency. Extensive experiments demonstrated the superiority of Fossil, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
Collectively, our framework redefines the design principles of efficient adaptation: Fossil provides theoretical grounding for dimension-wise decomposition, Fossil delivers practical efficiency parity with LoRA. This progression demonstrates that parameter efficiency need not come at the cost of expressivity, paving the way for adaptive fine-tuning in resource-constrained environments.
7 Limitations and Future work
Due to hardware constraints, we were unable to conduct and report results from the full-scale training that this method ideally requires. Nonetheless, through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of Fossil. The strong performance achieved on large language models gives us substantial confidence in the potential of this approach. We believe that Fossil can be readily applied to a wide range of multimodal tasks.
As a pioneering approach, Fossil still leaves several aspects open for deeper exploration. We hope that future research will conduct broader and more in-depth studies to further refine PEFT techniques and identify the most effective strategies for large language models.
Appendix A Appendix
A.1 RWKV7
Table 6: We fine-tuned LLMs using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
Model Strategy Trainable GSM8K Math HumanEval MT-Bench
RWKV7-3B Base 0M 44.35 - - -
LoRA 47.2M 55.64 - - -
PiSSA 47.2M 57.16 - -
DiSHA(Bone) 47.2M 58.22 - - -
Table 7: Hyperparameter settings for fine-tuning llama2-7B,Mistral-7B,RWKV6-7B,Qwen3-4B on NLG tasks
Hyperparameters LoRA DoRA PiSSA DiSHA(Bone)
Rank r 36 36 36 64
$\alpha$ 72 72 36 -
Dropout 0.0
Optimizer AdamW
LR 2e-5
LR Scheduler Cosine decay
Batch size 64
Warmup ratio 0.0
Epochs 1
Where Q,K,V,O,Up,Down,Gate
Table 8: Hyperparameter settings for fine-tuning llama2-13B on NLG tasks
Hyperparameters LoRA DoRA PiSSA DiSHA(Bone)
Rank r 64 64 64 128
$\alpha$ 128 128 64 -
Dropout 0.0
Optimizer AdamW
LR 2e-5
LR Scheduler Cosine decay
Batch size 128
Warmup ratio 0.0
Epochs 1
Where Q,K,V,O,Up,Down,Gate
Appendix B Settings of Experiments in No Free Lunch
Table 9: Experimental Setup: Datasets and Hyperparameters
General Configuration
Parameter Value
Random Seed (SEED) 43
Device (DEVICE) CUDA (if available, else CPU)
Base Model Architecture (MLP)
Input Dimension 64
Hidden Dimension 64
Output Dimension 64
Synthetic Dataset Generation
Base Function $\sin(2\pi x)$
Modified Function $\sin(2\pi x)+0.3\cos(3\pi x)$
Input $x$ Range $[-1,1]$
Training Samples ($N\_TRAIN$) 50
Validation Samples ($N\_VALID$) 100
Training Noise Std. Dev. (NOISE_STD) 0.05
Validation Noise Std. Dev. 0.0
Training Parameters
Base Model LR (BASE_LR) 0.001
Adaptation LR (ADAPT_LR) 0.001
Base Model Epochs (BASE_EPOCHS) 250
Adaptation Epochs (ADAPT_EPOCHS) 100
Evaluation Interval (EVAL_INTERVAL) 10
Adapter-Specific Ranks
LoRA Rank 2
VeRA Rank 64
BlockDiagonal Rank 4
PiSSA Rank 2
DoRA Rank 1
ProLoRA Rank 2
AdaLoRA Rank 2
MoS Rank 2
Note: Other adapter-specific hyperparameters (e.g., LoRA scale, VeRA d_init_val, DoRA lora_alpha, ProLoRA unshared_rank_u, MoS shard_dim_ratio) primarily use their default values as defined in the respective adapter class implementations or are derived based on the rank within benchmark functions. Refer to the provided Python code for their specific configurations during experiments.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope. In this work, we introduce a novel parameter-efficient fine-tuning (PEFT) method called Fissol.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Yes, we discussed the limitations in the last section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: There is no theoretical result or proof in this paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have provided the settings for our experiments.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have uploaded the code for this paper in the supplemental materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have provided the details in both experiment sections and appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All experiments were conducted multiple times, and the results were averaged.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide detailed hyperparameter settings in the appendix to ensure reproducibility.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Full compliance with ethical norms is ensured.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper does not involve such content and focuses on objective experimental results.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not cover these topics.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All aspects have been acknowledged.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This work does not involve any asset-related issues.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not cover this topic.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not cover this topic.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [No]
  79. Justification: This paper does not use any Large Language Models (LLMs).
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v10

@@
 Existing LoRA variants often struggle to strike an optimal balance between adaptability (model performance and convergence speed) and efficiency (computational overhead, memory usage, and initialization time).
-This paper introduces Fossil(Framework for Optimal Shard Sharing Integration in LoRA), a novel PEFT approach that addresses this trade-off through a simple shard-sharing mechanism.
-Fossil leverages the insight that a low-rank adaptation can be achieved by decomposing the weight matrix into multiple fragment matrices and utilizing a shared, trainable common fragment.
+This paper introduces MiSS(Matrix Shard Sharing ), a novel PEFT approach that addresses this trade-off through a simple shard-sharing mechanism.
+MiSS leverages the insight that a low-rank adaptation can be achieved by decomposing the weight matrix into multiple fragment matrices and utilizing a shared, trainable common fragment.
 This method constructs the low-rank update matrix through the replication of these shared, partitioned shards.
-We also propose a hardware-efficient and broadly applicable implementation for Fossil.
-Extensive experiments conducted on a range of tasks, alongside a systematic analysis of computational performance, demonstrate Fossil’s superiority.
-The results show that Fossil significantly outperforms standard LoRA and its prominent variants in both model performance metrics and computational efficiency, including initialization speed and training throughput.
-By effectively balancing expressive power and resource utilization, Fossil offers a compelling solution for efficiently adapting large-scale models.
+We also propose a hardware-efficient and broadly applicable implementation for MiSS.
+Extensive experiments conducted on a range of tasks, alongside a systematic analysis of computational performance, demonstrate MiSS’s superiority.
+The results show that MiSS significantly outperforms standard LoRA and its prominent variants in both model performance metrics and computational efficiency, including initialization speed and training throughput.
+By effectively balancing expressive power and resource utilization, MiSS offers a compelling solution for efficiently adapting large-scale models.
 [Uncaptioned image]https://github.com/huggingface/peft
 1 Introduction
-Fine-tuning Large Language Models (LLMs) [16, 17] is a prevalent methodology for adapting these models to specific downstream tasks.
+Fine-tuning Large Language Models (LLMs) [16, 17, 25] is a prevalent methodology for adapting these models to specific downstream tasks.
 However, the comprehensive fine-tuning of all parameters typically incurs substantial computational costs.
@@
 Considering the aforementioned trade-off, our investigation reveals that a surprisingly simple strategy—constructing a low-rank matrix through the replication of a shared, partitioned shard—can achieve a compelling balance across these criteria.
-This approach, which we introduce as Fossil,  exhibits notable strengths.
-In terms of adaptability, Fossil demonstrates superior approximation fidelity compared to LoRA and several of its variants.
+This approach, which we introduce as MiSS,  exhibits notable strengths.
+In terms of adaptability, MiSS demonstrates superior approximation fidelity compared to LoRA and several of its variants.
 Its relatively substantial gradient norm (as illustrated in Figure 4) promotes rapid convergence, a key characteristic of its high adaptability.
-Furthermore, concerning efficiency, Fossil benefits from rapid initialization; its hardware-friendly, repetition-based initialization facilitates swift processing, particularly when scaling to large models with substantial parameter counts.
-The formulation of Fossil can also be expressed in an efficient and mathematically equivalent form that exhibits favorable time and space complexity.
+Furthermore, concerning efficiency, MiSS benefits from rapid initialization; its hardware-friendly, repetition-based initialization facilitates swift processing, particularly when scaling to large models with substantial parameter counts.
+The formulation of MiSS can also be expressed in an efficient and mathematically equivalent form that exhibits favorable time and space complexity.
 Table 1: A variety of LoRA variants are listed, each with its specific update formulation and initialization strategy for the low-rank matrices.
@@
 MoS $y=\bm{W}_{0}x+\bm{B}^{s}\bm{A}^{s}x$ $\bm{A}^{\text{pub/pri}},\bm{B}^{\text{pub/pri}}\sim 0$
-Fossil(Ours) $y=\bm{W_{0}}x+\operatorname{expand}(\bm{D})x$ $\bm{D}\sim 0$
-Fossile(Ours) $y=\bm{W_{0}}x+\bm{D^{\top}S}$ $\bm{D}\sim 0$
+MiSS(Ours) $y=\bm{W_{0}}x+\operatorname{expand}(\bm{D})x$ $\bm{D}\sim 0$
+MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D^{\top}S}$ $\bm{D}\sim 0$
 We perform scaling experiments to verify our preliminary conclusions.
-Our extensive evaluation shows that Fossil demonstrates excellent performance on understanding and generation tasks, significantly surpassing other baseline method.
-Moreover, we present the training dynamics i.e., loss curves, which Fossil also showcase a significantly faster convergence rate compared to other baselines.
-In generation tasks, Fossil demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variants across the board.
+Our extensive evaluation shows that MiSS demonstrates excellent performance on understanding and generation tasks, significantly surpassing other baseline method.
+Moreover, we present the training dynamics i.e., loss curves, which MiSS also showcase a significantly faster convergence rate compared to other baselines.
+In generation tasks, MiSS demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variants across the board.
 This demonstrates the feasibility of the framework.
-Furthermore, we validate the efficiency of Fossil  which enables it to significantly outperform LoRA and its variants in both computational efficiency and memory usage.
+Furthermore, we validate the efficiency of MiSS  which enables it to significantly outperform LoRA and its variants in both computational efficiency and memory usage.
 Our contributions can be summarized as threefolds:
@@
   2. 2.
-    In Section 4, we proposed Fossil, an efficient and adaptable structure with shard sharing mechanism.
+    In Section 4, we proposed MiSS, an efficient and adaptable structure with shard sharing mechanism.
   3. 3.
@@
 (*x.shape[:-1], x.size(-1)//self.r, self.r)\\
-, dim=-2) @ self.fossil
+, dim=-2) @ self.miss
 return y
-Figure 1: Left. Schematic diagram of the forward computation and initialization methods in LoRA and Fossil. Right. PyTorch-style pseudocode illustrating the implementation of Fossil.
+Figure 1: Left. Schematic diagram of the forward computation and initialization methods in LoRA and MiSS. Right. PyTorch-style pseudocode illustrating the implementation of MiSS.
 2 Preliminaries and Related Works
@@
 We considered diverse methods  [^1] We have not included methods such as LoRA-GA [21] or LoRA+ [7] in our current analysis.
-While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider Fossil to be largely orthogonal to them.
-Consequently, the analytical techniques employed in their study may still offer valuable insights for Fossil.: (1) Full-parameter finetuning [13].
+While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider MiSS to be largely orthogonal to them.
+Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning [13].
 (2) LoRA [8].
@@
 A plausible explanation for this discrepancy is that such methods further decompose LoRA matrices into shared components which may inherently constrain their expressive power.
-Our method Fossil reaches a relatively advanced performance comparing to other variants.
+Our method MiSS reaches a relatively advanced performance comparing to other variants.
 Figure 2: No Free Launch Experiment. Left. The training loss curves of all methods. Middle. Initialization time w/ parameters. Right. Training time w/ parameters.
@@
 Results.
-The efficacy (See Figure 2) of Fossil is evident: its strategic combination of parameter sharing and an efficient computational design culminates in rapid, scalable performance across both initialization and training stages.
+The efficacy (See Figure 2) of MiSS is evident: its strategic combination of parameter sharing and an efficient computational design culminates in rapid, scalable performance across both initialization and training stages.
 In contrast, while techniques like PiSSA demonstrate commendable adaptability, as shown in prior experiments, their reliance on computationally intensive Singular Value Decomposition for initialization significantly hampers their overall speed.
 Other approaches, such as VeRA and AdaLoRA, offer efficient initialization and computation; however, as previously discussed, they often achieve this at the cost of comparatively reduced adaptability.
-4 Fossil: Shard Sharing for the Performance and Efficiency Tradeoff
+4 MiSS: Shard Sharing for the Performance and Efficiency Tradeoff
 4.1 Method Overview
@@
 For instance, if a matrix’s rows or shards are constructed by repeating a limited set of independent elements, its effective rank is often much smaller than its full dimensions.
-Based on this insight, we propose Fossil, which defines the weight update $\Delta\bm{W}$ as a large matrix generated from a small trainable matrix $\bm{D}$ through an expansion operation.
+Based on this insight, we propose MiSS, which defines the weight update $\Delta\bm{W}$ as a large matrix generated from a small trainable matrix $\bm{D}$ through an expansion operation.
 The updating of $\bm{W}$ and the forward pass can be expressed as:
@@
 The shards are vertically concatenated to match the dimensions of $\bm{W_{0}}$.
-4.2 Efficient Implementation of Fossil
+4.2 Efficient Implementation of MiSS
 The above formulation is effective in the initialization process, as it only needs to initialize a small $\bm{D}$.
 However, directly computing $\operatorname{expand}(\bm{D})x$ has a time complexity of $O(bldk)$ and memory complexity of $O(dk)$ , which can be computationally intensive.
-It is obvious that Fossil can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{r\times d}$, where $r$ is a tunable rank parameter.
+It is obvious that MiSS can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{r\times d}$, where $r$ is a tunable rank parameter.
 Instead of partitioning the output dimension $d$, we divide the input dimension $k$ into $r$ blocks, each of size $g=\lfloor k/r\rfloor$ (for simplicity, assume $k$ is divisible by $r$).
@@
 This structure avoids storing the $d\times k$ matrix explicitly, requiring only $\bm{D}\in\mathbb{R}^{r\times d}$, significantly reducing memory usage.
-The efficient implementation of Fossil relies on an innovative input aggregation mechanism, namely blockwise input summation.
+The efficient implementation of MiSS relies on an innovative input aggregation mechanism, namely blockwise input summation.
 We highlight its advantages through the following steps: (1) Input Partitioning and Aggregation: The aggregation exploits local redundancy in the input, preserving critical information while reducing the computational dimensionality.
@@
 Figure 3: Left.
-Fossil consumes less GPU memory and achieves higher throughput(# tokens per second).
+MiSS consumes less GPU memory and achieves higher throughput(# tokens per second).
 GPU memory usage refers to the amount of memory occupied by an RTX 4090 card.
@@
 5 Experiments
-In this section, we evaluate the performance of Fossil on various benchmark datasets.
+In this section, we evaluate the performance of MiSS on various benchmark datasets.
 Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model.
@@
 To ensure fair comparisons, we select methods officially included in the PEFT repository to avoid discrepancies caused by inconsistent tuning implementations.
-Table 2: The results of fine-tuning RoBERTa-base using Fossil and various LoRA variants were compared on a subset of the GLUE benchmark.
+Table 2: The results of fine-tuning RoBERTa-base using MiSS and various LoRA variants were compared on a subset of the GLUE benchmark.
 Method Trainable MNLI SST-2 CoLA QNLI MRPC
@@
 PiSSA 0.236% 85.72±0.40 93.64±0.13 67.28±0.59 91.40±0.54 88.11±0.24
-Fossil 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60
-Table 3: We conducted fine-tuning of large language models using Fossil and multiple LoRA variants, evaluating their performance on GSM8K, Math, HumanEval, and MBPP benchmarks.
+MiSS 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60
+Table 3: We conducted fine-tuning of large language models using MiSS and multiple LoRA variants, evaluating their performance on GSM8K, Math, HumanEval, and MBPP benchmarks.
 All reported results are averaged over three independent runs to ensure robustness.
@@
 PiSSA 89.9M 43.89 6.92 22.15 37.84
-Fossil 87.0M 48.16 8.58 23.63 36.81
+MiSS 87.0M 48.16 8.58 23.63 36.81
 RWKV 6-7B [15] LoRA 88.1M 38.13 6.06 - -
 PiSSA 88.1M 40.48 6.12 - -
-Fossil 88.1M 41.73 6.52 - -
+MiSS 88.1M 41.73 6.52 - -
 Mistral-7B [9] LoRA 94.4M 62.85 15.82 35.71 46.11
@@
 PiSSA 94.4M 67.01 18.13 41.28 51.37
-Fossil 87.0M 68.92 18.85 42.07 61.33
+MiSS 87.0M 68.92 18.85 42.07 61.33
 Llama2-13B [18] LoRA 250M 56.18 12.60 31.79 37.82
@@
 PiSSA 250M 66.64 13.82 33.57 46.03
-Fossil 255M 68.64 15.74 38.15 47.91
+MiSS 255M 68.64 15.74 38.15 47.91
 Qwen3-4B [24] LoRA 74.3M 84.38 15.20 73.27 78.32
@@
 PiSSA 74.3M 85.78 26.00 75.01 78.04
-Fossil 70.1M 85.52 34.82 74.48 78.05
+MiSS 70.1M 85.52 34.82 74.48 78.05
 5.1 Experiments on Natural Language Understanding
@@
 Each experiment is conducted with 3 different random seeds, and the average performance is reported.
-As shown in Table 2, Fossil demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and PiSSA.
+As shown in Table 2, MiSS demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and PiSSA.
 5.2 Experiment on Natural Language Generation
 Setup.
-To verify the generalizability of Fossil, we conducted more comprehensive experiments on LLM.
+To verify the generalizability of MiSS, we conducted more comprehensive experiments on LLM.
 we conducted 3 more task finetuning experiments on LLM: math and code.
@@
 The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used.
-It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing Fossil with LoRA.
-To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of Fossil to 64, ensuring that Fossil uses fewer parameters than LoRA to demonstrate its superiority.
+It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing MiSS with LoRA.
+To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of MiSS to 64, ensuring that MiSS uses fewer parameters than LoRA to demonstrate its superiority.
 Each experiment is conducted with 2 different random seeds, and the average performance is reported.
-32641282560612182430rankgradient normftLoRAPiSSAFossil
+32641282560612182430rankgradient normftLoRAPiSSAMiSS
 Figure 4: Left.
-The loss curve for LLaMA2-7B fine-tuned on MetaMathQA shows Fossil’s superior fitting ability across architectures and parameter settings, with a notably rapid loss reduction in the first 100 steps, underscoring its effectiveness.
+The loss curve for LLaMA2-7B fine-tuned on MetaMathQA shows MiSS’s superior fitting ability across architectures and parameter settings, with a notably rapid loss reduction in the first 100 steps, underscoring its effectiveness.
 Right.
@@
 Results.
-As shown in Table 6, Fossil consistently achieves state-of-the-art performance across multiple models and evaluation metrics.
-Specifically, on LLaMA2-7B, Fossil surpasses PiSSA on most metrics while using fewer trainable parameters than both LoRA and DoRA, showcasing its parameter efficiency and superior performance.
-On RWKV6-7B, where all methods share the same number of trainable parameters, Fossil still achieves the best overall performance, indicating its strong adaptation capability to non-transformer architectures.
-Across all evaluated models, including Mistral-7B, LLaMA2-13B, and Qwen3-4B, Fossil maintains its leading position on key benchmarks, often achieving the highest scores with significantly fewer parameters—further highlighting its effectiveness and scalability.
+As shown in Table 6, MiSS consistently achieves state-of-the-art performance across multiple models and evaluation metrics.
+Specifically, on LLaMA2-7B, MiSS surpasses PiSSA on most metrics while using fewer trainable parameters than both LoRA and DoRA, showcasing its parameter efficiency and superior performance.
+On RWKV6-7B, where all methods share the same number of trainable parameters, MiSS still achieves the best overall performance, indicating its strong adaptation capability to non-transformer architectures.
+Across all evaluated models, including Mistral-7B, LLaMA2-13B, and Qwen3-4B, MiSS maintains its leading position on key benchmarks, often achieving the highest scores with significantly fewer parameters—further highlighting its effectiveness and scalability.
 5.3 Effect of Rank $r$
-This subsection explores the upper limits of the Fossil structure by varying the rank $r$ in the Fossil matrix.
+This subsection explores the upper limits of the MiSS structure by varying the rank $r$ in the MiSS matrix.
 Comparative experiments were conducted by fine-tuning LLaMA2-7B  [^3] We use LLaMA2-7B instead of LLaMA3-8B because we found that the LLaMA3 series is over-optimized on math-related tasks.
@@
 The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases.
-Notably, when $r=16$, the Fossil structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark.
+Notably, when $r=16$, the MiSS structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark.
 However, its performance on the Math benchmark is only 3.73.
@@
 Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
-Table 4: Comparing different values of rank $(r)$ on LLaMA2-7B with Fossil.
+Table 4: Comparing different values of rank $(r)$ on LLaMA2-7B with MiSS.
 Model Rank Trainable GSM8K Math
@@
 LoRA $O(dr+rk)$ $O(blr(d+k))$
-Fossil $O(dr)$ $O(bldk)$
-Fossile $O(dr)$ $O(blr(d+\frac{k}{r}))$
+MiSS $O(dr)$ $O(bldk)$
+MiSSe $O(dr)$ $O(blr(d+\frac{k}{r}))$
 5.4 Gradient Norm Analysis
-To further investigate the adaptive capabilities of Fossil, we analyzed its initial gradient norms in comparison to Full Fine-tuning (FT), LoRA, and PiSSA.
+To further investigate the adaptive capabilities of MiSS, we analyzed its initial gradient norms in comparison to Full Fine-tuning (FT), LoRA, and PiSSA.
 The magnitude of initial gradient norms is often correlated with faster convergence and the ability of a model to effectively adapt during fine-tuning.
@@
 While PiSSA’s gradient norms increase with higher ranks, they generally remain significantly lower than those achieved by full fine-tuning.
-Notably, Fossil demonstrates substantially larger initial gradient norms compared to both LoRA and PiSSA across various ranks.
-More importantly, the gradient norm profile of Fossil closely approximates that of Full Fine-tuning.
-This proximity to FT’s gradient characteristics suggests that Fossil is capable of inducing more significant initial updates, potentially leading to a training dynamic more akin to full fine-tuning.
+Notably, MiSS demonstrates substantially larger initial gradient norms compared to both LoRA and PiSSA across various ranks.
+More importantly, the gradient norm profile of MiSS closely approximates that of Full Fine-tuning.
+This proximity to FT’s gradient characteristics suggests that MiSS is capable of inducing more significant initial updates, potentially leading to a training dynamic more akin to full fine-tuning.
 5.5 Resource and efficiency
-Table  5 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Fossil, and Bat on a single 4090 GPU.
+Table  5 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA and MiSSon a single 4090 GPU.
 The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512.
-The results show that Fossil has the highest computational efficiency, being nearly 10% faster than LoRA while also being more memory-efficient.
-However, Bat incurs significantly higher memory usage due to large intermediate values and is slower in comparison.
+The results show that MiSS has the highest computational efficiency, being nearly 10% faster than LoRA while also being more memory-efficient.
 At the end of the table, we provide the actual resource costs for fine-tuning RWKV6 on the MetaMathQA dataset using 4 NVIDIA 4090 GPUs, with checkpoint techniques applied.
@@
 GPU $\times$ 1 LoRA 55.1M 12074 MB 3.62 kt/s
-Fossil 55.1M 11052 MB 3.99 kt/s
+MiSS 55.1M 11052 MB 3.99 kt/s
 GPU $\times$ 4 LoRA 55.1M 4$\times$15328 MB 15.6 kt/s
-Fossil 55.1M 4$\times$15304 MB 16.0 kt/s
+MiSS 55.1M 4$\times$15304 MB 16.0 kt/s
 6 Conclusion
@@
 This approach significantly reduces resource overhead and opens up new directions for PEFT techniques.
-Furthermore, we present a hardware-optimized version of Fossil, achieving remarkable gains in computational efficiency.
-Extensive experiments demonstrated the superiority of Fossil, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
-Collectively, our framework redefines the design principles of efficient adaptation: Fossil provides theoretical grounding for dimension-wise decomposition, Fossil delivers practical efficiency parity with LoRA.
+Furthermore, we present a hardware-optimized version of MiSS, achieving remarkable gains in computational efficiency.
+Extensive experiments demonstrated the superiority of MiSS, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
+Collectively, our framework redefines the design principles of efficient adaptation: MiSS provides theoretical grounding for dimension-wise decomposition, MiSS delivers practical efficiency parity with LoRA.
 This progression demonstrates that parameter efficiency need not come at the cost of expressivity, paving the way for adaptive fine-tuning in resource-constrained environments.
@@
 Due to hardware constraints, we were unable to conduct and report results from the full-scale training that this method ideally requires.
-Nonetheless, through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of Fossil.
+Nonetheless, through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of MiSS.
 The strong performance achieved on large language models gives us substantial confidence in the potential of this approach.
-We believe that Fossil can be readily applied to a wide range of multimodal tasks.
-As a pioneering approach, Fossil still leaves several aspects open for deeper exploration.
+We believe that MiSS can be readily applied to a wide range of multimodal tasks.
+As a pioneering approach, MiSS still leaves several aspects open for deeper exploration.
 We hope that future research will conduct broader and more in-depth studies to further refine PEFT techniques and identify the most effective strategies for large language models.
@@
 A.1 RWKV7
-Table 6: We fine-tuned LLMs using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
+Table 6: We fine-tuned LLMs using MiSS and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
 Model Strategy Trainable GSM8K Math HumanEval MT-Bench
@@
 PiSSA 47.2M 57.16 - -
-DiSHA(Bone) 47.2M 58.22 - - -
+MiSS 47.2M 58.22 - - -
 Table 7: Hyperparameter settings for fine-tuning llama2-7B,Mistral-7B,RWKV6-7B,Qwen3-4B on NLG tasks
-Hyperparameters LoRA DoRA PiSSA DiSHA(Bone)
+Hyperparameters LoRA DoRA PiSSA MiSS
 Rank r 36 36 36 64
@@
 Table 8: Hyperparameter settings for fine-tuning llama2-13B on NLG tasks
-Hyperparameters LoRA DoRA PiSSA DiSHA(Bone)
+Hyperparameters LoRA DoRA PiSSA MiSS
 Rank r 64 64 64 128
@@
 VeRA Rank 64
-BlockDiagonal Rank 4
+MiSSRank 4
 PiSSA Rank 2
@@
 Refer to the provided Python code for their specific configurations during experiments.
-NeurIPS Paper Checklist
-The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
-Do not remove the checklist: The papers not including the checklist will be desk rejected.
-The checklist should follow the references and follow the (optional) supplemental material.
-The checklist does NOT count towards the page limit.
-Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
-  • You should answer [Yes] , [No] , or [N/A] .
-  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
-  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
-The checklist answers are an integral part of your paper submission.
-They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
-You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
-The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
-While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
-In general, answering "[No] " or "[N/A] " is not grounds for rejection.
-While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
-All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
-If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
-IMPORTANT, please:
-  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
-  • Keep the checklist subsection headings, questions/answers and guidelines below.
-  • Do not modify the questions and only use the provided macros for your answers.
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
-In this work, we introduce a novel parameter-efficient fine-tuning (PEFT) method called Fissol.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Yes, we discussed the limitations in the last section.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: There is no theoretical result or proof in this paper.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We have provided the settings for our experiments.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We have uploaded the code for this paper in the supplemental materials.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We have provided the details in both experiment sections and appendix.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: All experiments were conducted multiple times, and the results were averaged.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We provide detailed hyperparameter settings in the appendix to ensure reproducibility.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: Full compliance with ethical norms is ensured.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: This paper does not involve such content and focuses on objective experimental results.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This paper does not cover these topics.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: All aspects have been acknowledged.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: This work does not involve any asset-related issues.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: This paper does not cover this topic.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: This paper does not cover this topic.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [No]
-  79. Justification: This paper does not use any Large Language Models (LLMs).
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v11

@@
-Balancing LoRA Performance and Efficiency with Simple Shard Sharing
+MiSS: Revisiting the Trade-off in LoRA with an Efficient Shard-Sharing Structure
 Abstract
-Parameter-Efficient Fine-Tuning (PEFT) methods, particularly Low-Rank Adaptation (LoRA), effectively reduce the number of trainable parameters in Large Language Models (LLMs).
-However, as model scales continue to grow, the demand for computational resources remains a significant challenge.
-Existing LoRA variants often struggle to strike an optimal balance between adaptability (model performance and convergence speed) and efficiency (computational overhead, memory usage, and initialization time).
-This paper introduces MiSS(Matrix Shard Sharing ), a novel PEFT approach that addresses this trade-off through a simple shard-sharing mechanism.
-MiSS leverages the insight that a low-rank adaptation can be achieved by decomposing the weight matrix into multiple fragment matrices and utilizing a shared, trainable common fragment.
-This method constructs the low-rank update matrix through the replication of these shared, partitioned shards.
-We also propose a hardware-efficient and broadly applicable implementation for MiSS.
-Extensive experiments conducted on a range of tasks, alongside a systematic analysis of computational performance, demonstrate MiSS’s superiority.
-The results show that MiSS significantly outperforms standard LoRA and its prominent variants in both model performance metrics and computational efficiency, including initialization speed and training throughput.
-By effectively balancing expressive power and resource utilization, MiSS offers a compelling solution for efficiently adapting large-scale models.
+Low-Rank Adaptation (LoRA) is a widely adopted technique for parameter-efficient fine-tuning, but its slow convergence has spurred the development of numerous variants.
+Nevertheless, existing methods often fail to improve performance, memory footprint, and computational efficiency simultaneously.
+To address this challenge, we revisit the causes of LoRA’s slow convergence.
+Building on these insights, we propose Matrix Shard Sharing (MiSS), which updates shards of the original weight matrix using a single shared trainable matrix $\bm{D}$, initialized to zeros.
+To simultaneously ensure computational efficiency, low memory footprint, and scalable serving, we introduce MiSSe.
+Both theoretical analysis and empirical results demonstrate that our method reduces optimization complexity without compromising performance, thereby achieving a more favorable trade-off among performance, memory, and efficiency.
+Furthermore, we conduct a comprehensive comparative analysis of various PEFT methods, evaluating their memory usage, initialization overhead, and computational efficiency.
+By mapping the Pareto frontier across these dimensions, we show that MiSS occupies a favorable position, effectively capturing the advantages of prior approaches.
+https://github.com/Joluck/MiSS
 [Uncaptioned image]https://github.com/huggingface/peft
 1 Introduction
-Fine-tuning Large Language Models (LLMs) [16, 17, 25] is a prevalent methodology for adapting these models to specific downstream tasks.
-However, the comprehensive fine-tuning of all parameters typically incurs substantial computational costs.
-Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques [23] have been developed to mitigate the training expenditure associated with these large-scale models.
-Among such techniques, Low-Rank Adaptation (LoRA) [8] has distinguished itself as one of the most prominent PEFT methods.
+Fine-tuning Large Language Models (LLMs) (Radford et al., 2019; Raffel et al., 2020; Yin et al., 2024) is a prevalent methodology for adapting these models to specific downstream tasks.
+However, full fine-tuning of all parameters is computationally prohibitive.
+Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques (Xu et al., 2023) have been developed to mitigate the training expenditure associated with these large-scale models.
+Among such techniques, Low-Rank Adaptation (LoRA) (Hu et al., 2021) has distinguished itself as one of the most prominent PEFT methods.
 LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead.
@@
 Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
-The improvements of LoRA are mimicked in two streams: (1) Adaptability [6, 12, 2]: This refers to the celerity with which the method converges to an optimal or near-optimal state.
+Many improvements to LoRA have been proposed in recent years, which can be broadly categorized into two major streams: (1) Adaptability (Ding et al., 2023; Liu et al., 2024; Biderman et al., 2024): This refers to the convergence speed at which the method reaches an optimal or near-optimal state.
 The approximation must exhibit a representational capacity comparable to that of the original, full $\Delta\bm{W}$.
 Extensive experiments have shown that LoRA’s convergence is significantly slower compared to full fine-tuning.
-To address this issue, researchers have proposed several LoRA variants [7, 14, 20].
+To address this issue, researchers have proposed several LoRA variants (Hayou et al., 2024; Meng et al., 2024; Wang et al., 2024a).
 By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed.
 Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients.
-Certain solutions may improve the performance of finetuning, but makes the initialization process significantly more complex or increase the training cost, leading to a worse complexity.
-(2) Efficiency [11, 19, 22]: This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead.
+However, many of these methods overlook issues of computational efficiency and overall training overhead.
+For example, PiSSA (Meng et al., 2024) requires a lengthy initialization process, while LoRA-GA (Wang et al., 2024b) depends on modifications to the optimizer, resulting in incompatibility with certain optimizers.
+(2) Efficiency (Kopiczko et al., 2024; Wang et al., 2024c; 2025): This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead.
 Optimizing LoRA from an efficiency perspective can lead to reduced VRAM consumption and an accelerated training process.
-However, further simplification of the LoRA structure may result in an additional loss in performance.
-However, the critical balance between achieving highly expressive architectures and maintaining algorithmic efficiency has often received insufficient attention.
-In this work, we systematically investigate the Pareto frontier of the trade-off between Adaptability and Efficiency in PEFT.
-We contend that an ideal PEFT method must adeptly reconcile these two crucial aspects.
-To this end, we first conduct a series of foundational experiments focusing on these dimensions, encompassing a simulated pre-training and fine-tuning pipeline, computational complexity analysis, and initialization time tests.
-Our analysis reveals that while LoRA-like methods leveraging shard sharing can achieve excellent computational efficiency, their expressive power may be constrained by such sharing, sometimes leading to performance below the standard LoRA baseline.
-Conversely, SVD-based modifications to the original LoRA often exhibit strong adaptability, yet their practical efficiency can be undermined by complex initialization procedures and more intricate forward and update mechanisms.
-Considering the aforementioned trade-off, our investigation reveals that a surprisingly simple strategy—constructing a low-rank matrix through the replication of a shared, partitioned shard—can achieve a compelling balance across these criteria.
-This approach, which we introduce as MiSS,  exhibits notable strengths.
-In terms of adaptability, MiSS demonstrates superior approximation fidelity compared to LoRA and several of its variants.
-Its relatively substantial gradient norm (as illustrated in Figure 4) promotes rapid convergence, a key characteristic of its high adaptability.
-Furthermore, concerning efficiency, MiSS benefits from rapid initialization; its hardware-friendly, repetition-based initialization facilitates swift processing, particularly when scaling to large models with substantial parameter counts.
-The formulation of MiSS can also be expressed in an efficient and mathematically equivalent form that exhibits favorable time and space complexity.
+Although LoRA has demonstrated significant advantages in reducing parameter scale and computational cost, its effectiveness still falls short of fully matching full fine-tuning.
+To address this gap, researchers have proposed an increasing number of LoRA variants that gradually approach the performance of full fine-tuning.
+This raises a natural question:
+Given the inherent challenge for LoRA and its variants to balance performance, memory, and efficiency, how can we achieve an effective trade-off among all three dimensions?
 Table 1: A variety of LoRA variants are listed, each with its specific update formulation and initialization strategy for the low-rank matrices.
@@
 Method Forward Initialization
-LoRA [8] $y=\bm{W}_{0}x+\bm{BA}x$ $\bm{A}\sim N(0,\sigma^{2})$ $\bm{B}\sim 0$
+LoRA $y=\bm{W}_{0}x+\bm{BA}x$ $\bm{A}\sim N(0,\sigma^{2})$ $\bm{B}\sim 0$
 PiSSA $y=\bm{W}_{0}x+\bm{BA}x$ $\bm{A}=U_{[:,:r]}S_{[:r,:r]}^{1/2},\ \bm{B}=S_{[:r,:r]}^{1/2}V_{[:,:r]}^{\top}$
@@
 MiSS(Ours) $y=\bm{W_{0}}x+\operatorname{expand}(\bm{D})x$ $\bm{D}\sim 0$
-MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D^{\top}S}$ $\bm{D}\sim 0$
-We perform scaling experiments to verify our preliminary conclusions.
-Our extensive evaluation shows that MiSS demonstrates excellent performance on understanding and generation tasks, significantly surpassing other baseline method.
-Moreover, we present the training dynamics i.e., loss curves, which MiSS also showcase a significantly faster convergence rate compared to other baselines.
-In generation tasks, MiSS demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variants across the board.
-This demonstrates the feasibility of the framework.
-Furthermore, we validate the efficiency of MiSS  which enables it to significantly outperform LoRA and its variants in both computational efficiency and memory usage.
-Our contributions can be summarized as threefolds:
+MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D^{\intercal}S}$ $\bm{D}\sim 0$
+To strike a balance between performance, memory, and efficiency, we re-examined the key factors affecting LoRA’s slow convergence.
+Through an analysis of $S^{2}$FT (Yang et al., 2024) and LoRA+ (Hayou et al., 2024), we identified a critical phenomenon: During the LoRA fine-tuning process, both matrices $B$ and $A$ need to be updated simultaneously, which increases the complexity of optimization and ultimately leads to slower convergence.
+LoRA+ alleviates this issue by modifying the initial gradients, allowing the fine-tuning process to approximate full fine-tuning better.
+In contrast, $S^{2}$FT fixes one matrix as an orthogonal matrix, reducing the degrees of freedom in parameter updates and lowering optimization complexity, thereby enabling faster alignment with the optimal update direction.
+Inspired by these insights, we hypothesize that training only a single matrix could simplify optimization without sacrificing expressive capacity.
+We therefore propose Matrix Shard Sharing (MiSS), a method that updates a shard of the original weight matrix using a single, shared trainable matrix $\bm{D}$, initialized to zero.
+Thus, our approach maintains the low-rank property of the matrices while offering a more efficient alternative to $\bm{BA}$ updates in terms of computation.
+Gradient Norm Analysis.
+We analyze the initial gradient norm to verify our preliminary conclusions.
+In the experimental sections of the PiSSA, $S^{2}$FT, and LoRA-GA papers, we observed that LoRA exhibits a very small initial gradient norm compared to full fine-tuning, which shows a much larger one.
+Notably, all these improved methods share a common characteristic: their initial gradient norms are significantly larger than LoRA, and their early-stage convergence speed is comparable to that of full fine-tuning.
+Motivated by this, we evaluated the initial gradient norms of different methods across various models and datasets to examine whether MiSS follows the same pattern as other LoRA variants.
+The experimental results (Figure1,4) confirm that MiSS indeed shares this property, i.e., a larger initial gradient norm and faster early convergence.
+This also supports the hypothesis that optimizing a single matrix is inherently simpler.
+326412825602468Matrix rankGradientNorm
+326412825601234Matrix rankGradientNorm
+3264128256081624Matrix rankGradientNorm
+32641282560481216Matrix rankGradientNormFinetuneLoRAPiSSAMiSS
+Figure 1: Comparison of initial gradient norms across different training methods and the effect of rank. Results are shown for LLaMA2-7B and Qwen3-4B on the Math and Code datasets.
+MiSSe
+To achieve both computational efficiency and structural generality, we introduce an efficient and mathematically equivalent formulation of MiSS that offers improved time and space complexity.
+We further conduct extensive experiments (Table 5.3) to validate its effectiveness.
+We first evaluate MiSS on both Natural Language Understanding (NLU) and Generation (NLG) tasks, assessing its performance and scalability.
+Our results show that MiSS consistently outperforms LoRA and its variants across diverse LLM architectures, establishing new state-of-the-art results on a wide range of metrics.
+We then analyze the Pareto frontier of the adaptability-efficiency trade-off in PEFT.
+We argue that an ideal PEFT method should effectively balance these two essential dimensions.
+To this end, we conduct a series of foundational experiments, including a simulated pre-training and fine-tuning pipeline, computational complexity analysis, and initialization time evaluation.
+With comprehensive empirical results, we demonstrate that MiSS achieves a favorable balance across three key dimensions performance, memory, and efficiency, highlighting its practicality as a general PEFT solution.
+Our contributions can be summarized as follows:
   1. 1.
-    In Section 3, we discuss the balancing Between Adaptability and Efficiency in LoRA-like methods with highly artificial controlled dataset to make the preliminary verification transparently and easy for replication.
+    We propose MiSS, an efficient and adaptable structure with a shard-sharing mechanism, striking an effective balance among three essential properties—performance, memory efficiency, and computational efficiency.
   2. 2.
-    In Section 4, we proposed MiSS, an efficient and adaptable structure with shard sharing mechanism.
-  3. 3.
-    In Section 5, we scale the experiments across a wide range of datasets and models, validated its effectiveness through extensive experiments.
-def init(in_feats: int, out_feats: int):
-weight = nn.Parameter(
-torch.empty((out_features, in_features)) )
-D = nn.Parameter(torch.zeros(self.r, out_features))
-def forward(self, x):
-result = F.linear(x, self.weight) # x: [B, T, C]
-y = result + torch.sum(x.reshape
-(*x.shape[:-1], x.size(-1)//self.r, self.r)\\
-, dim=-2) @ self.miss
-return y
-Figure 1: Left. Schematic diagram of the forward computation and initialization methods in LoRA and MiSS. Right. PyTorch-style pseudocode illustrating the implementation of MiSS.
+    Through large-scale experiments across diverse datasets and model architectures, we provide a comprehensive evaluation of multiple PEFT methods.
+Our empirical results conclusively demonstrate that MiSS achieves a superior balance among these three properties compared to existing alternatives.
 2 Preliminaries and Related Works
@@
 One compelling venue is to change the form of LoRA.
-PiSSA [14] optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction.
-Using Singular Value Decomposition (SVD), OLoRA [3] leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal.
+PiSSA (Meng et al., 2024) optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction.
+Using Singular Value Decomposition (SVD), OLoRA (Büyükakyüz, 2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal.
 This orthogonal initialization helps maintain the stability of the parameter space during optimization.
@@
 LoRA+ extended this method by introducing independent learning rates for matrices $\bm{A}$ and $\bm{B}$ with a fixed ratio, improving the method’s efficiency.
-DoRA [12] decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately.
+DoRA (Liu et al., 2024) decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately.
 This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning.
@@
 Conversely, efficiency encompasses the method’s judicious use of computational resources, specifically time and memory.
-We utilize highly artificial controlled dataset and model with relatively small parameter count to make the verification transparently and easy for replication.
-We considered diverse methods  [^1] We have not included methods such as LoRA-GA [21] or LoRA+ [7] in our current analysis.
+We utilize highly artificial controlled dataset and model with a relatively small parameter count to make the verification transparently and easy for replication.
+We considered diverse methods  [^1] We have not included methods such as LoRA-GA (Wang et al., 2024b) or LoRA+ (Hayou et al., 2024) in our current analysis.
 While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider MiSS to be largely orthogonal to them.
-Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning [13].
-(2) LoRA [8].
-(3) Alternatives to LoRA w/ different architectures, including: PiSSA [14], VeRA [11], DoRA [12] and MoRA [10].
-(4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA [19], MoS [22].
+Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning (Lv et al., 2024).
+(2) LoRA (Hu et al., 2021).
+(3) Alternatives to LoRA w/ different architectures, including: PiSSA (Meng et al., 2024), VeRA (Kopiczko et al., 2024), DoRA (Liu et al., 2024) and MoRA (Jiang et al., 2024).
+(4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA (Wang et al., 2024c), MoS (Wang et al., 2025).
 (1) An overview of their forward form, initialization method can be found at Table 1.
@@
 Comprehensive details regarding the experimental settings are elaborated in Appendix B.
+Figure 2: No Free Launch Experiment. Left. The training loss curves of all methods. Middle. Initialization time w/ parameters. Right. Training time w/ parameters.
 Results.
@@
 Conversely, efficiency-focused techniques like MoS exhibit higher losses.
-A plausible explanation for this discrepancy is that such methods further decompose LoRA matrices into shared components which may inherently constrain their expressive power.
+A plausible explanation for this discrepancy is that such methods further decompose LoRA matrices into shared components, which may inherently constrain their expressive power.
 Our method MiSS reaches a relatively advanced performance comparing to other variants.
-Figure 2: No Free Launch Experiment. Left. The training loss curves of all methods. Middle. Initialization time w/ parameters. Right. Training time w/ parameters.
 3.2 Efficiency Analysis of LoRA Variants
@@
 4.1 Method Overview
-In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{AB}^{T}$, where $\bm{A}\in\mathbb{R}^{d\times r}$, $\bm{B}\in\mathbb{R}^{k\times r}$, and the rank $r\ll\min(d,k)$.
+In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{AB}^{\intercal}$, where $\bm{A}\in\mathbb{R}^{d\times r}$, $\bm{B}\in\mathbb{R}^{k\times r}$, and the rank $r\ll\min(d,k)$.
 This approach achieves efficiency by limiting the number of parameters.
@@
 For instance, if a matrix’s rows or shards are constructed by repeating a limited set of independent elements, its effective rank is often much smaller than its full dimensions.
+def init(in_features: int, in_features: int, rank: int):
+self.r = rank
+self.weight = $n$.Parameter(torch.empty((out_features, in_features)))
+self.D = $n$.Parameter(torch.zeros(self.r, out_features))
+def forward(self, x):
+result = F.linear(x, self.weight) # x: [B, T, C]
+y = result + x @ self.D.expand(in_features//self.r,1)
+return y
+Figure 3: Left. Structural diagram of $\Delta\bm{W}$ in LoRA and MiSS. Right. PyTorch-style pseudocode illustrating the implementation of MiSS.
 Based on this insight, we propose MiSS, which defines the weight update $\Delta\bm{W}$ as a large matrix generated from a small trainable matrix $\bm{D}$ through an expansion operation.
@@
 Formally:
-$(\operatorname{expand}(\bm{D}))^{\top}=\begin{bmatrix}(\mathbf{1}_{s_{1}}\bm{D_{1}})^{\top}\ (\mathbf{1}_{s_{2}}\bm{D_{2}})^{\top}\ \dots\ (\mathbf{1}_{s_{N}}\bm{D_{N}})^{\top}\end{bmatrix}$ (3)
+$(\operatorname{expand}(\bm{D}))^{\intercal}=\begin{bmatrix}(\mathbf{1}_{s_{1}}\bm{D_{1}})^{\intercal}\ (\mathbf{1}_{s_{2}}\bm{D_{2}})^{\intercal}\ \dots\ (\mathbf{1}_{s_{N}}\bm{D_{N}})^{\intercal}\end{bmatrix}$ (3)
 Here, $\bm{1}_{s_{i}}\in\mathbb{R}^{s_{i}\times 1}$ is an all-ones vector, and $\bm{1}_{s_{i}}\bm{D_{i}}$ denotes $\bm{D_{i}}$ repeated $s_{i}$ times vertically.
@@
 This enjoys the following updating term and forward pass:
-$\Delta\bm{Wx}=\bm{D^{\top}S},\ \ \bm{y}=\bm{W_{0}x}+\bm{D^{\top}S},\ \text{where}\ \bm{D}^{\top}\in\mathbb{R}^{d\times r}.$ (6)
-Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{D}^{\top}\bm{S}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
-This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=D^{T}S$.
+$\Delta\bm{Wx}=\bm{D^{\intercal}S},\ \ \bm{y}=\bm{W_{0}x}+\bm{D^{\intercal}S},\ \text{where}\ \bm{D}^{\intercal}\in\mathbb{R}^{d\times r}.$ (6)
+Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{D}^{\intercal}\bm{S}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
+This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=D^{\intercal}S$.
 Specifically, $\operatorname{expand}(\bm{D})\in\mathbb{R}^{d\times k}$ has rows corresponding to rows of $\bm{D}$, repeated across blocks in the $k$-dimension.
@@
 For example, with $k=1024$ and $r=16$, memory is reduced by about 64 times.
-An overall analysis of space and time complexity is analyzed in Table 5.
-(a) Loss-Token
-(b) Loss-Time
-Figure 3: Left.
-MiSS consumes less GPU memory and achieves higher throughput(# tokens per second).
-GPU memory usage refers to the amount of memory occupied by an RTX 4090 card.
-Right.
-The results of fine-tuning LLaMA2-7B on MetaMathQA.
+An overall analysis of space and time complexity is analyzed in Table 5.3.
 5 Experiments
 In this section, we evaluate the performance of MiSS on various benchmark datasets.
-Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model.
+Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset .
 Subsequently, we evaluated the Natural Language Generation (NLG [^2] Evaluations were performed with the OpenCompass repository, and the Math dataset was evaluated using a 5-shot prompt configuration.) capabilities by fine-tuning the LLM.
-To ensure fair comparisons, we select methods officially included in the PEFT repository to avoid discrepancies caused by inconsistent tuning implementations.
 Table 2: The results of fine-tuning RoBERTa-base using MiSS and various LoRA variants were compared on a subset of the GLUE benchmark.
@@
 MiSS 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60
-Table 3: We conducted fine-tuning of large language models using MiSS and multiple LoRA variants, evaluating their performance on GSM8K, Math, HumanEval, and MBPP benchmarks.
+Table 3: We conduct a systematic comparison of LoRA, DoRA, PiSSA, and MiSS across several mainstream large language models (Llama2, RWKV, Mistral, and Qwen3).
 All reported results are averaged over three independent runs to ensure robustness.
 The first-place entry should be highlighted in bold, and the second-place entry should be underlined.
-Model Strategy Trainable GSM8K [5] Math [26] HumanEval [4] Mbpp [1]
-Llama2-7B [18] LoRA 89.9M 40.75 5.22 17.74 35.15
+Model Strategy Trainable GSM8K Math HumanEval Mbpp
+Llama2-7B (Touvron et al., 2023) LoRA 89.9M 40.75 5.22 17.74 35.15
 DoRA 91.3M 42.93 6.51 21.95 36.53
@@
 MiSS 87.0M 48.16 8.58 23.63 36.81
-RWKV 6-7B [15] LoRA 88.1M 38.13 6.06 - -
+RWKV 6-7B (Peng et al., 2024) LoRA 88.1M 38.13 6.06 - -
 PiSSA 88.1M 40.48 6.12 - -
 MiSS 88.1M 41.73 6.52 - -
-Mistral-7B [9] LoRA 94.4M 62.85 15.82 35.71 46.11
+Mistral-7B (Jiang et al., 2023) LoRA 94.4M 62.85 15.82 35.71 46.11
 DoRA 95.8M 63.68 13.60 38.41 48.73
@@
 MiSS 87.0M 68.92 18.85 42.07 61.33
-Llama2-13B [18] LoRA 250M 56.18 12.60 31.79 37.82
+Llama2-13B (Touvron et al., 2023) LoRA 250M 56.18 12.60 31.79 37.82
 DoRA 252M 61.56 13.60 33.50 39.25
@@
 MiSS 255M 68.64 15.74 38.15 47.91
-Qwen3-4B [24] LoRA 74.3M 84.38 15.20 73.27 78.32
+Qwen3-4B (Yang et al., 2025) LoRA 74.3M 84.38 15.20 73.27 78.32
 DoRA 75.4M 85.11 21.73 74.20 78.77
@@
 we conducted 3 more task finetuning experiments on LLM: math and code.
-(1) Math: We trained our model on a 395k subset of MetaMathQA [26], a dataset bootstrapped from other math instruction tuning datasets like GSM8K [5] and MATH  [26], with higher complexity and diversity.
-(2) Code: We train our model on a 100k subset of CodeFeedback [27], a high-quality code instruction dataset, removing explanations after code blocks.
-The model is tested on HumanEval [4] and Mbpp [1].
+(1) Math: We trained our model on a 395k subset of MetaMathQA (Yu et al., 2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K (Cobbe et al., 2021) and MATH  (Yu et al., 2023), with higher complexity and diversity.
+(2) Code: We train our model on a 100k subset of CodeFeedback (Zheng et al., 2024), a high-quality code instruction dataset, removing explanations after code blocks.
+The model is tested on HumanEval (Chen et al., 2021) and Mbpp (Austin et al., 2021).
 The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used.
@@
 Each experiment is conducted with 2 different random seeds, and the average performance is reported.
-32641282560612182430rankgradient normftLoRAPiSSAMiSS
-Figure 4: Left.
-The loss curve for LLaMA2-7B fine-tuned on MetaMathQA shows MiSS’s superior fitting ability across architectures and parameter settings, with a notably rapid loss reduction in the first 100 steps, underscoring its effectiveness.
-Right.
-The initial gradient norm of model training.
 Results.
-As shown in Table 6, MiSS consistently achieves state-of-the-art performance across multiple models and evaluation metrics.
-Specifically, on LLaMA2-7B, MiSS surpasses PiSSA on most metrics while using fewer trainable parameters than both LoRA and DoRA, showcasing its parameter efficiency and superior performance.
-On RWKV6-7B, where all methods share the same number of trainable parameters, MiSS still achieves the best overall performance, indicating its strong adaptation capability to non-transformer architectures.
-Across all evaluated models, including Mistral-7B, LLaMA2-13B, and Qwen3-4B, MiSS maintains its leading position on key benchmarks, often achieving the highest scores with significantly fewer parameters—further highlighting its effectiveness and scalability.
+(a) Loss-Token
+(b) Loss-Time
+Figure 4: Loss curves of LLaMA2-7B fine-tuned on MetaMathQA using LoRA and MiSS(̇a) Loss vs. tokens. (b) Loss vs. training time.
+The results (See Table  3) show that MiSS consistently achieves the best or near-best performance across different model scales and diverse tasks.
+For example, on Llama2-7B, MiSS significantly surpasses LoRA on GSM8K (48.16 vs.
+40.75) and also improves HumanEval and Mbpp with scores of 8.58 and 36.81, respectively.
+As the model scales up to Llama2-13B, MiSS maintains its advantage, particularly on code generation benchmarks such as HumanEval and Mbpp, outperforming both PiSSA and DoRA.
+More importantly, on the latest large-scale model Qwen3-4B, which is trained on richer data, MiSS achieves the highest Math score of 34.82, far exceeding other methods, while remaining competitive on GSM8K and Mbpp.
+These findings demonstrate that MiSS is not only effective on medium-sized models but also scales robustly to larger architectures and data-rich models.
 5.3 Effect of Rank $r$
-This subsection explores the upper limits of the MiSS structure by varying the rank $r$ in the MiSS matrix.
-Comparative experiments were conducted by fine-tuning LLaMA2-7B  [^3] We use LLaMA2-7B instead of LLaMA3-8B because we found that the LLaMA3 series is over-optimized on math-related tasks.
-on the MetaMathQA dataset and validating on GSM8K and Math benchmarks.
-The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases.
-Notably, when $r=16$, the MiSS structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark.
-However, its performance on the Math benchmark is only 3.73.
-The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of $r$ impacts the model’s ability to understand unseen data.
-Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
+We evaluate MiSS with varying matrix ranks to study the trade-off between tuning capacity and parameter cost.
+The Table 4 reports results for ranks $r\in\{16,32,64,128\}$ (corresponding to $\{21.7\text{M},43.5\text{M},87.0\text{M},174.0\text{M}\}$ trainable parameters).
+Performance on GSM8K and the Math benchmark improves monotonically as the rank increases: GSM8K rises from 45.90 at $r=16$ to 53.49 at $r=128$, while Math increases from 3.77 to 10.08.
+In practice, $r=64$ offers a favorable trade-off (48.16 GSM8K, 8.58 Math) between performance gains and parameter overhead.
 Table 4: Comparing different values of rank $(r)$ on LLaMA2-7B with MiSS.
@@
 128 174.0M 53.49 10.08
-Figure 5: Time and space complexity. e denotes efficient form.
+Table 5: Time and space complexity. e denotes efficient form.
 Method Space Time
@@
 MiSSe $O(dr)$ $O(blr(d+\frac{k}{r}))$
-5.4 Gradient Norm Analysis
-To further investigate the adaptive capabilities of MiSS, we analyzed its initial gradient norms in comparison to Full Fine-tuning (FT), LoRA, and PiSSA.
-The magnitude of initial gradient norms is often correlated with faster convergence and the ability of a model to effectively adapt during fine-tuning.
-Our comparative analysis, illustrated in Figure 4 (Right), reveals distinct behaviors among the evaluated methods.
-Standard LoRA consistently exhibits the lowest initial gradient norms across the tested ranks.
-While PiSSA’s gradient norms increase with higher ranks, they generally remain significantly lower than those achieved by full fine-tuning.
-Notably, MiSS demonstrates substantially larger initial gradient norms compared to both LoRA and PiSSA across various ranks.
-More importantly, the gradient norm profile of MiSS closely approximates that of Full Fine-tuning.
-This proximity to FT’s gradient characteristics suggests that MiSS is capable of inducing more significant initial updates, potentially leading to a training dynamic more akin to full fine-tuning.
-5.5 Resource and efficiency
-Table  5 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA and MiSSon a single 4090 GPU.
-The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512.
-The results show that MiSS has the highest computational efficiency, being nearly 10% faster than LoRA while also being more memory-efficient.
-At the end of the table, we provide the actual resource costs for fine-tuning RWKV6 on the MetaMathQA dataset using 4 NVIDIA 4090 GPUs, with checkpoint techniques applied.
-Table 5: Resource and efficiency
-Devices Strategy Trainable Memory throughput
-GPU $\times$ 1 LoRA 55.1M 12074 MB 3.62 kt/s
-MiSS 55.1M 11052 MB 3.99 kt/s
-GPU $\times$ 4 LoRA 55.1M 4$\times$15328 MB 15.6 kt/s
-MiSS 55.1M 4$\times$15304 MB 16.0 kt/s
+5.4 Pareto plot
+The emergence of PEFT techniques is motivated by dual objectives: mitigating GPU memory constraints and exploring more efficient model architectures.
+Nevertheless, numerous contemporary studies disproportionately focus on ultimate performance benchmarks, overlooking critical practical considerations like computational efficiency and training duration—an emphasis that clearly diverges from the original rationale for PEFT.
+In this section, we undertake a multi-dimensional investigation into the relationships among computational overhead, efficiency, and performance for diverse models.
+Leveraging the official Hugging Face PEFT (Mangrulkar et al., 2022) benchmarking framework, our evaluations are conducted under fair and reproducible conditions.
+The Pareto frontiers in our evaluation provide definitive evidence of MiSS’s effectiveness.
+In every experimental setting, MiSS is uniquely positioned in the top-left corner—the optimal region—signifying that it delivers the best performance with minimal efficiency cost.
+This consistent advantage underscores MiSS’s unique contribution in balancing these competing objectives.
+Figure 5: Pareto front of MiSS comparing with other PEFT methods. We select three more methods as the baseline on the balancing of memory and performance.
+Table 6: Experimental results across PEFT methods on Llama-3.2-3B.
+PEFT Type Total Time Train Time Test Accuracy Train Loss Accelerator Memory (Bytes)
+Max Reserved 99th Reserved Avg
+RSLORA 2069 1871 0.5299 0.5657 22,538,092,544 17,953,927,987 12,128,059,444
+C3A 2125 1924 0.5102 0.5808 22,280,142,848 17,825,917,829 11,804,454,210
+[…]
+RANDLORA 2457 2213 0.5072 0.5785 22,798,139,392 18,436,063,232 12,743,670,025
+SHIRA 2085 1867 0.5072 0.5789 21,743,271,936 17,637,383,864 12,240,924,809
+OFT 2494 2214 0.5057 0.5947 22,294,822,912 17,939,310,837 12,057,354,384
+LORA 1993 1796 0.4822 0.6069 22,273,851,392 17,710,763,212 11,868,689,976
+DORA 2287 2023 0.4807 0.6068 24,553,455,616 19,189,150,515 12,490,471,636
+LORAFA 2026 1821 0.4299 0.6510 20,187,185,152 16,257,394,933 11,106,307,276
+LOHA 2591 2341 0.4185 0.6570 23,886,561,280 19,247,870,771 13,446,820,344
+IA3 1922 1746 0.4124 0.6569 23,135,780,864 18,398,356,439 12,023,331,867
+ADALORA 2209 1986 0.3904 0.6863 22,793,945,088 18,203,426,160 12,361,399,900
+LOKR 2352 2152 0.3753 0.6877 23,565,697,024 18,987,698,094 13,173,683,073
+P_TUNING 1918 1707 0.3707 0.6740 20,937,965,568 17,215,688,540 11,867,101,593
+VBLORA 2210 1962 0.3700 0.7143 22,181,576,704 17,635,223,797 11,735,344,663
+VERA 2025 1820 0.3685 0.6927 21,596,471,296 17,291,123,097 11,489,715,316
+BOFT 11,114 8292 0.3647 0.7268 24,427,626,496 20,103,445,872 14,814,855,089
+IA3 2005 1783 0.3450 0.7657 23,137,878,016 18,398,566,154 12,023,227,429
+TRAINABLE_TOKENS 1814 1572 0.2881 0.7862 20,956,839,936 16,957,675,929 12,730,137,942
+PROMPT_TUNING 2715 2394 0.2525 0.7790 24,408,752,128 20,650,676,715 15,297,364,466
+ADAPTION_PROMPT 2261 1989 0.2206 0.8317 22,410,166,272 17,907,664,814 11,893,757,234
+PREFIX_TUNING 1959 1662 0.1471 0.7887 20,912,799,744 16,945,051,074 11,766,684,083
+FOURIERFT 2824 2422 0.1198 0.9979 23,681,040,384 19,054,869,872 13,111,221,498
+PROMPT_TUNING 2700 2380 0.0500 1.0655 24,379,392,000 20,669,781,770 15,297,773,830
+FOURIERFT 2824 2424 0.0008 1.2480 23,653,777,408 19,017,267,937 13,104,129,350
+LN_TUNING 1870 1657 0.0000 1.2370 21,177,040,896 16,903,066,091 11,385,589,622
 6 Conclusion
-In this work, we proposed the Foissl framework, which divides pre-trained weights into multiple shards and updates them using a shared trainable matrix.
-This approach significantly reduces resource overhead and opens up new directions for PEFT techniques.
-Furthermore, we present a hardware-optimized version of MiSS, achieving remarkable gains in computational efficiency.
-Extensive experiments demonstrated the superiority of MiSS, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
-Collectively, our framework redefines the design principles of efficient adaptation: MiSS provides theoretical grounding for dimension-wise decomposition, MiSS delivers practical efficiency parity with LoRA.
-This progression demonstrates that parameter efficiency need not come at the cost of expressivity, paving the way for adaptive fine-tuning in resource-constrained environments.
+This work tackles the critical inefficiency of simultaneous matrix updates in Low-Rank Adaptation (LoRA), which leads to slow convergence and suboptimal resource use.
+We propose MiSS as a compelling solution—a new PEFT framework that updates decomposed weight shards using a single, shared matrix.
+This approach drastically reduces optimization complexity and resource demands.
+Comprehensive experiments validate that MiSS consistently outperforms existing methods in accuracy, memory footprint, and computational speed, offering a fundamentally more efficient pathway for adapting large models.
 7 Limitations and Future Work
-Due to hardware constraints, we were unable to conduct and report results from the full-scale training that this method ideally requires.
-Nonetheless, through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of MiSS.
+Through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of MiSS.
 The strong performance achieved on large language models gives us substantial confidence in the potential of this approach.
@@
 Modified Function $\sin(2\pi x)+0.3\cos(3\pi x)$
-Input $x$ Range $[-1,1]$
-Training Samples ($N\_TRAIN$) 50
-Validation Samples ($N\_VALID$) 100
+Input $x$ Range [1,1][-1,1]
+Training Samples (N_TRAINN\_TRAIN) 50
+Validation Samples (N_VALIDN\_VALID) 100
 Training Noise Std. Dev. (NOISE_STD) 0.05

### v12

@@
 1 Introduction
-Fine-tuning Large Language Models (LLMs) (Radford et al., 2019; Raffel et al., 2020; Yin et al., 2024) is a prevalent methodology for adapting these models to specific downstream tasks.
+Fine-tuning Large Language Models (LLMs) (radford2019language; raffel2020exploring; yin2024stablemask) is a prevalent methodology for adapting these models to specific downstream tasks.
 However, full fine-tuning of all parameters is computationally prohibitive.
-Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques (Xu et al., 2023) have been developed to mitigate the training expenditure associated with these large-scale models.
-Among such techniques, Low-Rank Adaptation (LoRA) (Hu et al., 2021) has distinguished itself as one of the most prominent PEFT methods.
+Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques (xu2023parameter) have been developed to mitigate the training expenditure associated with these large-scale models.
+Among such techniques, Low-Rank Adaptation (LoRA) (hu2021lora) has distinguished itself as one of the most prominent PEFT methods.
 LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead.
@@
 Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
-Many improvements to LoRA have been proposed in recent years, which can be broadly categorized into two major streams: (1) Adaptability (Ding et al., 2023; Liu et al., 2024; Biderman et al., 2024): This refers to the convergence speed at which the method reaches an optimal or near-optimal state.
+Many improvements to LoRA have been proposed in recent years, which can be broadly categorized into two major streams: (1) Adaptability (ding2023parameter; liu2024dora; biderman2024lora): This refers to the convergence speed at which the method reaches an optimal or near-optimal state.
 The approximation must exhibit a representational capacity comparable to that of the original, full $\Delta\bm{W}$.
 Extensive experiments have shown that LoRA’s convergence is significantly slower compared to full fine-tuning.
-To address this issue, researchers have proposed several LoRA variants (Hayou et al., 2024; Meng et al., 2024; Wang et al., 2024a).
+To address this issue, researchers have proposed several LoRA variants (hayou2024lora+; meng2024pissa; wang2024lora).
 By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed.
@@
 However, many of these methods overlook issues of computational efficiency and overall training overhead.
-For example, PiSSA (Meng et al., 2024) requires a lengthy initialization process, while LoRA-GA (Wang et al., 2024b) depends on modifications to the optimizer, resulting in incompatibility with certain optimizers.
-(2) Efficiency (Kopiczko et al., 2024; Wang et al., 2024c; 2025): This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead.
+For example, PiSSA (meng2024pissa) requires a lengthy initialization process, while LoRA-GA (wang2024loragalowrankadaptationgradient) depends on modifications to the optimizer, resulting in incompatibility with certain optimizers.
+(2) Efficiency (kopiczko2024veravectorbasedrandommatrix; wang2024prolorapartialrotationempowers; wang2025mosunleashingparameterefficiency): This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead.
 Optimizing LoRA from an efficiency perspective can lead to reduced VRAM consumption and an accelerated training process.
@@
 MiSS(Ours) $y=\bm{W_{0}}x+\operatorname{expand}(\bm{D})x$ $\bm{D}\sim 0$
-MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D^{\intercal}S}$ $\bm{D}\sim 0$
+MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D}\sum_{i=1}^{g}\bm{x^{(g)}}$ $\bm{D}\sim 0$
 To strike a balance between performance, memory, and efficiency, we re-examined the key factors affecting LoRA’s slow convergence.
-Through an analysis of $S^{2}$FT (Yang et al., 2024) and LoRA+ (Hayou et al., 2024), we identified a critical phenomenon: During the LoRA fine-tuning process, both matrices $B$ and $A$ need to be updated simultaneously, which increases the complexity of optimization and ultimately leads to slower convergence.
+Through an analysis of $S^{2}$FT (yang2024s2ftefficientscalablegeneralizable), LoRA-FA (zhang2023lorafamemoryefficientlowrankadaptation), and LoRA+ (hayou2024lora+), we identified a critical phenomenon:
+During the LoRA fine-tuning process, both matrices $B$ and AA need to be updated simultaneously, which increases the complexity of optimization and ultimately leads to slower convergence.
 LoRA+ alleviates this issue by modifying the initial gradients, allowing the fine-tuning process to approximate full fine-tuning better.
@@
 Motivated by this, we evaluated the initial gradient norms of different methods across various models and datasets to examine whether MiSS follows the same pattern as other LoRA variants.
-The experimental results (Figure1,4) confirm that MiSS indeed shares this property, i.e., a larger initial gradient norm and faster early convergence.
+The experimental results (Figure1) confirm that MiSS indeed shares this property, i.e., a larger initial gradient norm and faster early convergence.
 This also supports the hypothesis that optimizing a single matrix is inherently simpler.
@@
 Figure 1: Comparison of initial gradient norms across different training methods and the effect of rank. Results are shown for LLaMA2-7B and Qwen3-4B on the Math and Code datasets.
-MiSSe
-To achieve both computational efficiency and structural generality, we introduce an efficient and mathematically equivalent formulation of MiSS that offers improved time and space complexity.
-We further conduct extensive experiments (Table 5.3) to validate its effectiveness.
+Efficient Implementation
+To achieve better computational efficiency, we introduce MiSSe, an alternative design that maintains the core principle of parameter sharing while offering improved time and space complexity through input-dimension aggregation.
+We further conduct extensive experiments (Table 2) to validate its effectiveness.
 We first evaluate MiSS on both Natural Language Understanding (NLU) and Generation (NLG) tasks, assessing its performance and scalability.
@@
 One compelling venue is to change the form of LoRA.
-PiSSA (Meng et al., 2024) optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction.
-Using Singular Value Decomposition (SVD), OLoRA (Büyükakyüz, 2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal.
+PiSSA (meng2024pissa) optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction.
+Using Singular Value Decomposition (SVD), OLoRA (buyukakyuz2024olora) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal.
 This orthogonal initialization helps maintain the stability of the parameter space during optimization.
@@
 LoRA+ extended this method by introducing independent learning rates for matrices $\bm{A}$ and $\bm{B}$ with a fixed ratio, improving the method’s efficiency.
-DoRA (Liu et al., 2024) decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately.
+DoRA (liu2024dora) decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately.
 This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning.
@@
 However, they also introduce more complex initialization steps and increase preprocessing time.
-3 No Free Launch: Balancing Between Adaptability and Efficiency
+3 No Free Lunch: Balancing Between Adaptability and Efficiency
 This section elucidates the fundamental trade-off inherent in LoRA-style PEFT techniques: the delicate balance between their adaptability and efficiency.
@@
 We utilize highly artificial controlled dataset and model with a relatively small parameter count to make the verification transparently and easy for replication.
-We considered diverse methods  [^1] We have not included methods such as LoRA-GA (Wang et al., 2024b) or LoRA+ (Hayou et al., 2024) in our current analysis.
+We considered diverse methods  [^1] We have not included methods such as LoRA-GA (wang2024loragalowrankadaptationgradient) or LoRA+ (hayou2024lora+) in our current analysis.
 While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider MiSS to be largely orthogonal to them.
-Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning (Lv et al., 2024).
-(2) LoRA (Hu et al., 2021).
-(3) Alternatives to LoRA w/ different architectures, including: PiSSA (Meng et al., 2024), VeRA (Kopiczko et al., 2024), DoRA (Liu et al., 2024) and MoRA (Jiang et al., 2024).
-(4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA (Wang et al., 2024c), MoS (Wang et al., 2025).
+Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning (lv2024parameterfinetuninglargelanguage).
+(2) LoRA (hu2021lora).
+(3) Alternatives to LoRA w/ different architectures, including: PiSSA (meng2024pissa), VeRA (kopiczko2024veravectorbasedrandommatrix), DoRA (liu2024dora) and MoRA (jiang2024morahighrankupdatingparameterefficient).
+(4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA (wang2024prolorapartialrotationempowers), MoS (wang2025mosunleashingparameterefficiency).
 (1) An overview of their forward form, initialization method can be found at Table 1.
@@
 Following this pre-training, the target function is subtly altered, and an additional dataset sampled from this modified function is employed for training to assess the adaptation performance of various fine-tuning techniques.
-Comprehensive details regarding the experimental settings are elaborated in Appendix B.
+Comprehensive details regarding the experimental settings are elaborated in Appendix C.
 Figure 2: No Free Launch Experiment. Left. The training loss curves of all methods. Middle. Initialization time w/ parameters. Right. Training time w/ parameters.
@@
 4.1 Method Overview
-In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{AB}^{\intercal}$, where $\bm{A}\in\mathbb{R}^{d\times r}$, $\bm{B}\in\mathbb{R}^{k\times r}$, and the rank $r\ll\min(d,k)$.
+In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{BA}$, where $\bm{A}\in\mathbb{R}^{r\times k}$, $\bm{B}\in\mathbb{R}^{d\times r}$, and the rank $r\ll\min(d,k)$.
 This approach achieves efficiency by limiting the number of parameters.
@@
 However, directly computing $\operatorname{expand}(\bm{D})x$ has a time complexity of $O(bldk)$ and memory complexity of $O(dk)$, which can be computationally intensive.
-It is obvious that MiSS can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{r\times d}$, where $r$ is a tunable rank parameter.
+It is obvious that MiSS can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{d\times r}$, where $r$ is a tunable rank parameter.
 Instead of partitioning the output dimension $d$, we divide the input dimension $k$ into $r$ blocks, each of size $g=\lfloor k/r\rfloor$ (for simplicity, assume $k$ is divisible by $r$).
 For an input $\bm{x}\in\mathbb{R}^{b\times l\times k}$, partition it along the $k$-dimension, and sum each block along the $k$-dimension:
-$\bm{x}$ $=[\bm{x^{(1)}},\bm{x^{(2)}},\dots,\bm{x^{(r)}}],\quad\bm{x^{(i)}}\in\mathbb{R}^{b\times l\times g},$ (4)
-$\bm{S}$ $=\left[\sum_{j=1}^{g}\bm{x^{(1)}}_{[:,:,j]},\sum_{j=1}^{g}\bm{x^{(2)}}_{[:,:,j]},\dots,\sum_{j=1}^{g}\bm{x^{(r)}}_{[:,:,j]}\right]\in\mathbb{R}^{b\times l\times r}.$ (5)
+$\bm{x^{(i)}}$ $=\bm{x_{[:,:,(i-1)*r:i*r]}}\in\mathbb{R}^{b\times l\times r}$ (4)
+$\bm{x}$ $=[\bm{x^{(1)}},\bm{x^{(2)}},\dots,\bm{x^{(g)}}]$ (5)
+$\bm{S}$ $=\sum_{i=1}^{g}\bm{x^{(g)}}\in\mathbb{R}^{b\times l\times r}$ (6)
 This enjoys the following updating term and forward pass:
-$\Delta\bm{Wx}=\bm{D^{\intercal}S},\ \ \bm{y}=\bm{W_{0}x}+\bm{D^{\intercal}S},\ \text{where}\ \bm{D}^{\intercal}\in\mathbb{R}^{d\times r}.$ (6)
-Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{D}^{\intercal}\bm{S}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
-This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=D^{\intercal}S$.
+$\Delta\bm{Wx}=\bm{DS},\ \ \bm{y}=\bm{W_{0}x}+\bm{DS},\ \text{where}\ \bm{D}\in\mathbb{R}^{d\times r}.$ (7)
+Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{DS}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
+This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=\bm{DS}$.
 Specifically, $\operatorname{expand}(\bm{D})\in\mathbb{R}^{d\times k}$ has rows corresponding to rows of $\bm{D}$, repeated across blocks in the $k$-dimension.
 E.g., if $k=6$, $r=3$, and $g=2$, the $i$-th row of $\operatorname{expand}(\bm{D})$ takes values $\bm{D}_{j,i}$ in block $j=\lceil j^{\prime}/g\rceil$, where $j^{\prime}$ is the column index.
-This structure avoids storing the $d\times k$ matrix explicitly, requiring only $\bm{D}\in\mathbb{R}^{r\times d}$, significantly reducing memory usage.
+This structure avoids storing the $d\times k$ matrix explicitly, requiring only $\bm{D}\in\mathbb{R}^{d\times r}$, significantly reducing memory usage.
 The efficient implementation of MiSS relies on an innovative input aggregation mechanism, namely blockwise input summation.
@@
 (3) Resource Savings: Memory usage drops comparing to original form.
-For example, with $k=1024$ and $r=16$, memory is reduced by about 64 times.
-An overall analysis of space and time complexity is analyzed in Table 5.3.
+4.3 Systematic Analysis of Memory and Efficiency for LoRA and MiSS
+This subsection systematically compares LoRA variants against MiSS, dissecting their intrinsic differences in memory consumption (governed by parameter count) and computational efficiency (governed by FLOPs and operator type).
+Our analysis centers on the core update formulations: $\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}$ for LoRA, versus $\Delta\mathbf{W}\mathbf{x}=\mathbf{D}\mathbf{S}$ for the efficient form of MiSS (MiSSe), where $\mathbf{S}$ denotes the blockwise input aggregation.
+We denote the LoRA rank as $r_{\text{L}}$, MiSS rank as $r_{\text{M}}$, with input dimension $k$ and output dimension $d$.
+Limitations of LoRA Variants: Parameter Reduction $\neq$ Computational Speedup
+As illustrated in Table 2, there exists a fundamental misalignment between parameter efficiency and computational cost in existing PEFT methods.
+While variants like AdaLoRA, DoRA, and VeRA significantly reduce Trainable Parameters (TPs) through novel initialization or decomposition strategies, they almost universally inherit the sequential matrix multiplication logic $\mathbf{B}(\mathbf{A}\mathbf{x})$.
+Consequently, their Space Complexity and FLOPs remain bound by the $O((d+k)\times r)$ lower limit.
+Furthermore, sophisticated variants such as LoHA introduce additional structural overhead (e.g., the $2r$ factor), causing actual memory occupancy and latency to exceed the original LoRA despite having fewer trainable parameters.
+Table 2: Comparison of PEFT Methods.
+Note that while distinct LoRA variants reduce TPs, they fail to improve Space Complexity and FLOPs due to the unchanged sequential computation, unlike the proposed MiSS.
+Methods Space Complexity FLOPs TPs
+FT $O(d\times k)$ $O(d\times k)$ $d\cdot k$
+LoRA $O((d+k)\times r)$ $O((d+k)\times r)$ $(d+k)\cdot r$
+LoRA-FA $O((d+k)\times r)$ $O((d+k)\times r)$ $d\cdot r$
+AdaLoRA $\times r) O((d+k+r)×r)O((d+k$$\times r)$ $(d+k)\cdot r+r^{2}$
+LoHA $O(2r\times(d+k))$ $O(2r\times(d+k))$ $2\cdot(d+k)\cdot r$
+VeRA $O((d+k)r+r+d)$ $O((d+k)r+r+d)$ $d+r$
+MiSSe $\bm{O(d\times r)}$ $\bm{O(k+d\times r)}$ $\bm{d\cdot r}$
+Single-Matrix Paradigm and Computational Decomposition
+MiSS fundamentally diverges from the standard LoRA architecture by employing a single low-rank matrix $\mathbf{D}\in\mathbb{R}^{r_{1}\times r_{2}}$ , rather than the dual-matrix structure ($\mathbf{A},\mathbf{B}$).
+Crucially, we observe that $\mathbf{D}$ in MiSSe is dimensionally consistent with $\mathbf{B}$ in LoRA, as both correspond to the output dimension $d$ and function as the output operation matrix.
+This structural alignment allows us to naturally decompose the computation into two distinct stages: Input Transformation ($C_{\text{Step 1}}$) and Output Projection ($C_{\text{Step 2}}$).
+This insight isolates the efficiency distinction entirely to $C_{\text{Step 1}}$.
+While LoRA relies on an expensive matrix multiplication ($\mathbf{A}\mathbf{x}$), MiSSe utilizes a cost-efficient block summation ($\operatorname{sum}(\mathbf{x})$).
+The comparative analysis is summarized below:
+Table 3: Computational Decomposition of MiSSe vs. LoRA
+Metric LoRA MiSSe
+Structure Dual Matrices ($\mathbf{A},\mathbf{B}$) Single Matrix ($\mathbf{D}$)
+$C_{\text{Step 2}}$ (Output Projection) Matrix Mult. $\mathbf{B}\mathbf{h}$ ($d\times r$) Matrix Mult. $\mathbf{D}\mathbf{S}$ ($d\times r$)
+$C_{\text{Step 1}}$ (Input Transform) Matrix Mult. $\mathbf{A}\mathbf{x}$ ($O(BLkr)$) Block Sum $\operatorname{sum}(\mathbf{x})$ ($O(BLk)$)
+Parameter Count ($N$) $O(r(k+d))$ $O(rd)$
+Total FLOPs $O(BL(kr+rd))$ $O(BL(k+rd))$
 5 Experiments
-In this section, we evaluate the performance of MiSS on various benchmark datasets.
-Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset .
-Subsequently, we evaluated the Natural Language Generation (NLG [^2] Evaluations were performed with the OpenCompass repository, and the Math dataset was evaluated using a 5-shot prompt configuration.) capabilities by fine-tuning the LLM.
+In this section, we conduct a comprehensive set of experiments to validate the effectiveness and generalizability of MiSS across diverse domains.
+We assess performance on a wide range of tasks, including language, image, and video benchmarks.
+Specifically, we evaluate Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset, and Natural Language Generation (NLG) capabilities by fine-tuning various large language models (LLMs).
+We extend our evaluation to multimodal settings using the VTAB-1K benchmark to demonstrate the robust adaptability of MiSS beyond textual domains.
+Furthermore, we provide a detailed analysis of the Pareto frontier (Section 5.3) to definitively illustrate MiSS’s superior computational efficiency and minimal hardware overhead when compared to existing Parameter-Efficient Fine-Tuning (PEFT) methods.
+5.1 Superior Performance across Language and Vision Domains
+MiSS demonstrates exceptional versatility, maintaining a commanding lead or highly competitive performance across diverse benchmarks in both the language and vision domains. (Setup B)
+Natural Language Understanding (NLU).
+On the GLUE benchmark (Table 4), fine-tuning RoBERTa-base with MiSS showcases notable strength.
+It achieves an outstanding result on the challenging CoLA dataset (72.86), significantly surpassing LoRA and PiSSA.
+This performance indicates superior data-fitting capabilities and faster convergence on complex linguistic tasks.
 Table 4: The results of fine-tuning RoBERTa-base using MiSS and various LoRA variants were compared on a subset of the GLUE benchmark.
-Method Trainable MNLI SST-2 CoLA QNLI MRPC
-LoRA 0.236% 85.63±0.01 94.03±0.02 62.40±0.71 91.37±0.97 87.98±0.23
-PiSSA 0.236% 85.72±0.40 93.64±0.13 67.28±0.59 91.40±0.54 88.11±0.24
-MiSS 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60
+Method Trainable MNLI SST-2 CoLA QNLI MRPC Avg
+LoRA 0.236% 85.63±0.01 94.03±0.02 62.40±0.71 91.37±0.97 87.98±0.23 84.28
+PiSSA 0.236% 85.72±0.40 93.64±0.13 67.28±0.59 91.40±0.54 88.11±0.24 85.23
+MiSS 0.236% 85.71±0.32 93.60±0.07 72.86±3.13 91.43±0.76 88.14±0.60 86.35
+Natural Language Generation (NLG).
+Across five mainstream LLMs (Llama2, Mistral, RWKV, Qwen3), MiSS consistently achieves the best or near-best average performance (Table 5).
+Notably, it demonstrates substantial gains in complex reasoning tasks, recording the highest Math score (34.82) on Qwen3-4B and the highest average score (47.79) on Mistral-7B.
+These findings highlight that MiSS is not only effective on medium-sized models but also scales robustly to larger architectures and data-rich models.
 Table 5: We conduct a systematic comparison of LoRA, DoRA, PiSSA, and MiSS across several mainstream large language models (Llama2, RWKV, Mistral, and Qwen3).
@@
 The first-place entry should be highlighted in bold, and the second-place entry should be underlined.
-Model Strategy Trainable GSM8K Math HumanEval Mbpp
-Llama2-7B (Touvron et al., 2023) LoRA 89.9M 40.75 5.22 17.74 35.15
-DoRA 91.3M 42.93 6.51 21.95 36.53
-PiSSA 89.9M 43.89 6.92 22.15 37.84
-MiSS 87.0M 48.16 8.58 23.63 36.81
-RWKV 6-7B (Peng et al., 2024) LoRA 88.1M 38.13 6.06 - -
-PiSSA 88.1M 40.48 6.12 - -
-MiSS 88.1M 41.73 6.52 - -
-Mistral-7B (Jiang et al., 2023) LoRA 94.4M 62.85 15.82 35.71 46.11
-DoRA 95.8M 63.68 13.60 38.41 48.73
-PiSSA 94.4M 67.01 18.13 41.28 51.37
-MiSS 87.0M 68.92 18.85 42.07 61.33
-Llama2-13B (Touvron et al., 2023) LoRA 250M 56.18 12.60 31.79 37.82
-DoRA 252M 61.56 13.60 33.50 39.25
-PiSSA 250M 66.64 13.82 33.57 46.03
-MiSS 255M 68.64 15.74 38.15 47.91
-Qwen3-4B (Yang et al., 2025) LoRA 74.3M 84.38 15.20 73.27 78.32
-DoRA 75.4M 85.11 21.73 74.20 78.77
-PiSSA 74.3M 85.78 26.00 75.01 78.04
-MiSS 70.1M 85.52 34.82 74.48 78.05
-5.1 Experiments on Natural Language Understanding
-We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC.
-Performance is evaluated on the development set using accuracy as the primary metric.
-The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU.
-Each experiment is conducted with 3 different random seeds, and the average performance is reported.
-As shown in Table 2, MiSS demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and PiSSA.
-5.2 Experiment on Natural Language Generation
-Setup.
-To verify the generalizability of MiSS, we conducted more comprehensive experiments on LLM.
-we conducted 3 more task finetuning experiments on LLM: math and code.
-(1) Math: We trained our model on a 395k subset of MetaMathQA (Yu et al., 2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K (Cobbe et al., 2021) and MATH  (Yu et al., 2023), with higher complexity and diversity.
-(2) Code: We train our model on a 100k subset of CodeFeedback (Zheng et al., 2024), a high-quality code instruction dataset, removing explanations after code blocks.
-The model is tested on HumanEval (Chen et al., 2021) and Mbpp (Austin et al., 2021).
-The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used.
-It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing MiSS with LoRA.
-To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of MiSS to 64, ensuring that MiSS uses fewer parameters than LoRA to demonstrate its superiority.
-Each experiment is conducted with 2 different random seeds, and the average performance is reported.
-Results.
-(a) Loss-Token
-(b) Loss-Time
-Figure 4: Loss curves of LLaMA2-7B fine-tuned on MetaMathQA using LoRA and MiSS(̇a) Loss vs. tokens. (b) Loss vs. training time.
-The results (See Table  3) show that MiSS consistently achieves the best or near-best performance across different model scales and diverse tasks.
-For example, on Llama2-7B, MiSS significantly surpasses LoRA on GSM8K (48.16 vs.
-40.75) and also improves HumanEval and Mbpp with scores of 8.58 and 36.81, respectively.
-As the model scales up to Llama2-13B, MiSS maintains its advantage, particularly on code generation benchmarks such as HumanEval and Mbpp, outperforming both PiSSA and DoRA.
-More importantly, on the latest large-scale model Qwen3-4B, which is trained on richer data, MiSS achieves the highest Math score of 34.82, far exceeding other methods, while remaining competitive on GSM8K and Mbpp.
-These findings demonstrate that MiSS is not only effective on medium-sized models but also scales robustly to larger architectures and data-rich models.
+Model Strategy Trainable GSM8K Math HumanEval Mbpp Avg
+Llama2-7B (touvron2023llama) LoRA 89.9M 40.75 5.22 17.74 35.15 24.72
+DoRA 91.3M 42.93 6.51 21.95 36.53 26.48
+PiSSA 89.9M 43.89 6.92 22.15 37.84 27.70
+MiSS 87.0M 48.16 8.58 23.63 36.81 29.30
+RWKV 6-7B (peng2024eagle) LoRA 88.1M 38.13 6.06 - - 22.10
+PiSSA 88.1M 40.48 6.12 - - 23.30
+MiSS 88.1M 41.73 6.52 - - 24.13
+Mistral-7B (jiang2023mistral7b) LoRA 94.4M 62.85 15.82 35.71 46.11 40.12
+DoRA 95.8M 63.68 13.60 38.41 48.73 41.10
+PiSSA 94.4M 67.01 18.13 41.28 51.37 44.45
+MiSS 87.0M 68.92 18.85 42.07 61.33 47.79
+Llama2-13B (touvron2023llama) LoRA 250M 56.18 12.60 31.79 37.82 34.60
+DoRA 252M 61.56 13.60 33.50 39.25 36.98
+PiSSA 250M 66.64 13.82 33.57 46.03 39.52
+MiSS 255M 68.64 15.74 38.15 47.91 42.11
+Qwen3-4B (yang2025qwen3technicalreport) LoRA 74.3M 84.38 15.20 73.27 78.32 62.79
+DoRA 75.4M 85.11 21.73 74.20 78.77 64.95
+PiSSA 74.3M 85.78 26.00 75.01 78.04 66.21
+MiSS 70.1M 85.52 34.82 74.48 78.05 68.22
+Vision Task
+To validate the ability of MiSS to adapt to non-textual tasks, we conducted experiments on the VTAB-1K image and video benchmarks (Table 6).
+MiSS achieved an average accuracy of 88.02 on image tasks and 72.96 on video tasks, making it highly competitive with top-performing baseline methods like LoRA and DoRA.
+Crucially, this competitive performance is delivered with a significantly lower parameter budget ($\approx 0.4$ #TPs) compared to LoRA/DoRA ($\approx 0.8$ #TPs), confirming that the efficiency of MiSS transcends the language domain and is applicable to multimodal foundation models.
+Table 6: Performance comparison on VTAB-1K image and video benchmarks.Results are adopted from SliceFine (kowsher2025slicefineuniversalwinningslicehypothesis).
+Method Image Video
+Caltech Flowers Pets Camel. Euro. Retino. KITTI Avg #TPs UCF101 Kinetics HMDB Avg #TPs
+Full 89.92 97.41 85.87 81.65 88.12 73.62 77.93 84.93 85.83 92.30 55.23 65.79 74.99 86.65
+VeRA 91.53 99.19 91.04 86.45 92.97 74.25 77.92 87.62 0.240 92.28 57.21 66.77 72.09 0.242
+LoRA 92.03 99.18 90.92 87.73 92.65 74.23 80.42 88.08 0.833 93.88 57.81 67.37 73.02 0.835
+DoRA 91.86 99.27 91.08 85.88 91.42 75.28 80.46 87.89 0.834 92.84 57.77 67.33 72.65 0.836
+MiSS 92.14 99.23 91.05 86.28 92.83 73.71 80.91 88.02 0.414 93.82 57.75 67.31 72.96 0.415
 5.2 Effect of Rank $r$
-We evaluate MiSS with varying matrix ranks to study the trade-off between tuning capacity and parameter cost.
-The Table 4 reports results for ranks $r\in\{16,32,64,128\}$ (corresponding to $\{21.7\text{M},43.5\text{M},87.0\text{M},174.0\text{M}\}$ trainable parameters).
-Performance on GSM8K and the Math benchmark improves monotonically as the rank increases: GSM8K rises from 45.90 at $r=16$ to 53.49 at $r=128$, while Math increases from 3.77 to 10.08.
-In practice, $r=64$ offers a favorable trade-off (48.16 GSM8K, 8.58 Math) between performance gains and parameter overhead.
 Table 7: Comparing different values of rank $(r)$ on LLaMA2-7B with MiSS.
@@
 128 174.0M 53.49 10.08
-Table 5: Time and space complexity. e denotes efficient form.
-Method Space Time
-Full $O(dk)$ $O(bld(d+k))$
-LoRA $O(dr+rk)$ $O(blr(d+k))$
-MiSS $O(dr)$ $O(bldk)$
-MiSSe $O(dr)$ $O(blr(d+\frac{k}{r}))$
-5.4 Pareto plot
+We evaluate MiSS with varying matrix ranks to study the trade-off between tuning capacity and parameter cost.
+The Table 7 reports results for ranks $r\in\{16,32,64,128\}$ (corresponding to $\{21.7\text{M},43.5\text{M},87.0\text{M},174.0\text{M}\}$ trainable parameters).
+Performance on GSM8K and the Math benchmark improves monotonically as the rank increases: GSM8K rises from 45.90 at $r=16$ to 53.49 at $r=128$, while Math increases from 3.77 to 10.08.
+In practice, $r=64$ offers a favorable trade-off (48.16 GSM8K, 8.58 Math) between performance gains and parameter overhead.
+5.3 MiSS’s Superior Balance on the Pareto Frontier: Optimally Trading Off Efficiency and Performance
 The emergence of PEFT techniques is motivated by dual objectives: mitigating GPU memory constraints and exploring more efficient model architectures.
@@
 In this section, we undertake a multi-dimensional investigation into the relationships among computational overhead, efficiency, and performance for diverse models.
-Leveraging the official Hugging Face PEFT (Mangrulkar et al., 2022) benchmarking framework, our evaluations are conducted under fair and reproducible conditions.
+Leveraging the official Hugging Face PEFT (peft) benchmarking framework, our evaluations are conducted under fair and reproducible conditions.
 The Pareto frontiers in our evaluation provide definitive evidence of MiSS’s effectiveness.
@@
 7 Limitations and Future Work
-Through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of MiSS.
-The strong performance achieved on large language models gives us substantial confidence in the potential of this approach.
-We believe that MiSS can be readily applied to a wide range of multimodal tasks.
 As a pioneering approach, MiSS still leaves several aspects open for deeper exploration.
@@
 Appendix A Appendix
-A.1 RWKV7
+A.1 Additional Experiments
+(a) Loss-Token
+(b) Loss-Time
+Figure 5: Loss curves of LLaMA2-7B fine-tuned on MetaMathQA using LoRA and MiSS(̇a) Loss vs. tokens. (b) Loss vs. training time.
+A.2 RWKV7
 Table 9: We fine-tuned LLMs using MiSS and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
@@
 Where Q,K,V,O,Up,Down,Gate
-Appendix B Settings of Experiments in No Free Lunch
+Appendix B Settings of Experiments
+NLU
+We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC.
+Performance is evaluated on the development set using accuracy as the primary metric.
+The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU.
+Each experiment is conducted with 3 different random seeds, and the average performance is reported.
+As shown in Table 4, MiSS demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and PiSSA.
+NLG
+To verify the generalizability of MiSS, we conducted more comprehensive experiments on LLM.
+we conducted 3 more task finetuning experiments on LLM: math and code.
+(1) Math: We trained our model on a 395k subset of MetaMathQA (yu2023metamath), a dataset bootstrapped from other math instruction tuning datasets like GSM8K (cobbe2021gsm8k) and MATH  (yu2023metamath), with higher complexity and diversity.
+(2) Code: We train our model on a 100k subset of CodeFeedback (zheng2024opencodeinterpreter), a high-quality code instruction dataset, removing explanations after code blocks.
+The model is tested on HumanEval (chen2021evaluating) and Mbpp (austin2021programsynthesislargelanguage).
+The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used.
+It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing MiSS with LoRA.
+To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of MiSS to 64, ensuring that MiSS uses fewer parameters than LoRA to demonstrate its superiority.
+Each experiment is conducted with 2 different random seeds, and the average performance is reported.
+Vision Task
+on VTAB-1K image classification using ViT-Base-Patch16-224
+Appendix C Settings of Experiments in No Free Lunch
 Table 12: Experimental Setup: Datasets and Hyperparameters